### v3

# Estimates for trilinear and quadrilinear character sums

Abstract.
We obtain new bounds on some trilinear and quadrilinear character sums, which are non-trivial starting from very short ranges of the variables. An application to an apparently new problem on oscillations of characters on differences between Farey fractions is given. Other applications include a modular analogue of a multiplicative hybrid problem of Iwaniec and Sárközy (1987) and the solvability of some prime type equations with constraints.
Key words and phrases:
character sums, trilinear forms, quadrilinear forms
2020 Mathematics Subject Classification:
11L40
1. Introduction and backgrounds
1.1. Set-up
Motivated by various applications to analytic number theory, estimates for character sums received a lot of attention in the past decades. In many situations, the variables are supported over some additively structured sets, such as sets of consecutive integers. But the difficulty can vary significantly since the weights might be arbitrary. This paper aims to study two kinds of multilinear character sums with arbitrary weights.
Throughout this paper, denote by $p$ a large prime, and by $\chi$ a non-trivial multiplicative character modulo $p$. Take two integers $a,b$ with $p\nmid ab$. For $K,L,M,N\geqslant 1$, we consider the trilinear character sum
(1.1) $\mathfrak{T}(\bm{\alpha},\bm{\beta})=\sum_{k\leqslant K}\sum_{m\leqslant M}\sum_{n\leqslant N}\alpha_{m}\beta_{k,n}\chi(ak+bmn)$
and the quadrilinear sum
(1.2) $\mathfrak{Q}(\bm{\alpha},\bm{\beta})=\sum_{k\leqslant K}\sum_{\ell\leqslant L}\sum_{m\leqslant M}\sum_{n\leqslant N}\alpha_{\ell,m}\beta_{k,n}\chi(ak\ell+bmn),$
where $\bm{\alpha}=(\alpha_{m})$ or $(\alpha_{\ell,m})$ and $\bm{\beta}=(\beta_{k,n})$ are some complex weights. Note that one obtains (1.1) if taking $L=1$ and $\alpha_{1,m}=\alpha_{m}$ in (1.2). No confusion on the definitions of $\bm{\alpha}$ arises since one is for the trilinear sum, and the other is for the quadrilinear one.
In various practical applications, one aims to obtain upper bounds, as strong as possible, for $\mathfrak{T}(\bm{\alpha},\bm{\beta})$ and $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$ in terms of the $\ell_{\rho}$-norms $\|\bm{\alpha}\|_{\rho}$ and $\|\bm{\beta}\|_{\rho}$ (see (1.7) below for the definition of norms). We refer to the following inequalities
$|\mathfrak{T}(\bm{\alpha},\bm{\beta})|\leqslant\min\{\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|_{\infty}KMN,\leavevmode\nobreak\ \|\bm{\alpha}\|_{2}\|\bm{\beta}\|_{2}(KMN)^{\frac{1}{2}}\}$
and
$|\mathfrak{Q}(\bm{\alpha},\bm{\beta})|\leqslant\min\{\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|_{\infty}KLMN,\leavevmode\nobreak\ \|\bm{\alpha}\|_{2}\|\bm{\beta}\|_{2}(KLMN)^{\frac{1}{2}}\}$
as trivial bounds, which follow directly from the triangle inequality and the Cauchy–Schwarz inequality. Our aim is to beat the above trivial bounds for $\bm{\alpha},\bm{\beta}$ as general as possible and $K,L,M,N$ as small as possible compared to $p$.
We would like to emphasise that weights in $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$ depend on variables from different products, which makes treatments of such sums much more difficult, as the standard smoothing technique does not immediately apply.
Our bounds for $\mathfrak{T}(\bm{\alpha},\bm{\beta})$ and $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$ with power-savings, as well as applications, will be given in Section 2, and we would like to give a brief outline of related results right now.
1.2. Related bilinear sums
The above studies on character sums over sumsets can be dated back to Vinogradov [Vi54, Chapter V, Exercise 8.c] on the following bilinear form
(1.3) $\mathfrak{B}(\bm{\alpha},\bm{\beta})=\sum_{m\in{\mathcal{M}}}\sum_{n\in{\mathcal{N}}}\alpha_{m}\beta_{n}\chi(m+n),$
where $\chi$ is a non-trivial Dirichlet character modulo $p$ and ${\mathcal{M}},{\mathcal{N}}\subseteq[1,p]$ are arbitrary subsets with $M=\#{\mathcal{M}}$ and $N=\#{\mathcal{N}}$. A direct application of Fourier techniques yields
(1.4) $|\mathfrak{B}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{2}\|\bm{\beta}\|_{2}\leavevmode\nobreak\ p^{\frac{1}{2}}.$
Although this bound is widely known, its full derivation is not easy to find in the literature, however it can be found as a short proof of [SS17, Equation (1.4)]. Despite a very short and elementary proof, the bound (1.4) has never been improved in full generality. However, Karatsuba [Ka92] (see also [Ka93, Chapter VIII, Problem 9]) has proved, that if
(1.5) $M>p^{\frac{1}{2}+\eta}\qquad\mbox{and}\qquad N>p^{\eta}$
for some $\eta>0$, then the inequality
(1.6) $|\mathfrak{B}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|_{\infty}MNp^{-\kappa}$
holds with some $\kappa>0$, depending only on $\eta$. The range (1.5) reveals the Pólya–Vinogradov threshold even when summing over arbitrary subsets. A similar phenomenon can also be found in [Xi23]. We note that the proof of (1.6), in the range (1.5), does not seem to be present in literature, and a concise proof with an explicit upper bound, is provided in Appendix A to convince cautious readers.
It is worthwhile to point out that $\mathfrak{B}(\bm{\alpha},\bm{\beta})$ has received considerable attention in recent years due to its connection with the Paley graph conjecture. It is conjectured, for instance with just constant weights $\alpha_{m}=\beta_{n}=1$, that
$\mathfrak{B}(\bm{\alpha},\bm{\beta})=o(MN)$
for all subsets ${\mathcal{M}},{\mathcal{N}}$ with $M,N>p^{\eta}$ and $\eta>0$. This is far from proven, and the classical inequalities of Vinogradov (1.4) and Karatsuba (1.6) still stand. See [DBSW21, HP20] for recent progress towards this conjecture.
For structural sets, such as intervals, or sets with small doubling, or sets supported on short intervals, a large variety of bounds for $\mathfrak{B}(\bm{\alpha},\bm{\beta})$ are known. For example, the above bounds of Vinogradov and Karatsuba have been improved by Friedlander and Iwaniec [FI93, Theorem 3] when ${\mathcal{M}}$ and ${\mathcal{N}}$ are contained in intervals of lengths at most $\sqrt{p}$ and are of cardinalities $M,N\geqslant p^{\frac{11}{24}+\eta}$. Of course, the main point here is that $\frac{11}{24}<\frac{1}{2}$, which neither (1.4) nor (1.6) with (1.5) can achieve. The exponent $\frac{11}{24}$ has been reduced to $\frac{9}{20}$ by Bourgain, Garaev, Konyagin and Shparlinski [BGKS13, Theorem 25]. Several more results of this type can be found in [BGKS12, BKS15, Ch08, Ha17, Ka08, SS22, SS18, SV17, Vo18].
1.3. Related trilinear and quadrilinear sums
Before concluding this section, we also mention some recent works on various variants of $\mathfrak{B}(\bm{\alpha},\bm{\beta})$. To proceed, we assume ${\mathcal{H}},{\mathcal{K}},{\mathcal{M}},{\mathcal{N}}\subseteq[1,p]$ are arbitrary subsets. Hanson [Ha17] has studied
$\sum_{k\in{\mathcal{K}}}\sum_{m\in{\mathcal{M}}}\sum_{n\in{\mathcal{N}}}\alpha_{m}\beta_{n}\chi(k+m+n),$
while Roche-Newton, Shparlinski and Winterhof[RSW19] have considered
$\sum_{k\in{\mathcal{K}}}\sum_{m\in{\mathcal{M}}}\sum_{n\in{\mathcal{N}}}\chi(km+mn+nk)\mathrm{e}\left(\frac{a(km+mn+nk)}{p}\right)$
with $\gcd(a,p)=1$. Shkredov and Shparlinski [SS17] have treated quadrilinear forms
$\sum_{h\in{\mathcal{H}}}\sum_{k\in{\mathcal{K}}}\sum_{m\in{\mathcal{M}}}\sum_{n\in{\mathcal{N}}}\alpha_{h}\beta_{k,m,n}\chi(h+k+mn),$
$\sum_{h\in{\mathcal{H}}}\sum_{k\in{\mathcal{K}}}\sum_{m\in{\mathcal{M}}}\sum_{n\in{\mathcal{N}}}\alpha_{h}\beta_{k,m,n}\chi(h+k(m+n)).$
See also[KMPS20] for some recent generalizations and refinements.
Our main object $\mathfrak{T}(\bm{\alpha},\bm{\beta})$ is intimately related to the following trilinear character sum
$\sum_{k\leqslant K}\sum_{m\in{\mathcal{M}}}\sum_{n\in{\mathcal{N}}}\alpha_{k}\beta_{m}\gamma_{n}\chi(k+mn),$
where $\chi$ is a non-trivial Dirichlet character modulo $p$ and ${\mathcal{M}},{\mathcal{N}}\subseteq[1,p]$ are arbitrary subsets with $M=\#{\mathcal{M}}$ and $N=\#{\mathcal{N}}$. Banks and Shparlinski [BS20, Theorem 2.2] give the upper bound
$KMN(p^{-1}+(KM)^{-1}+K^{-2})^{\frac{1}{2r}}(p^{\frac{1}{4r}}+N^{-\frac{1}{2}}p^{\frac{1}{2r}})p^{o(1)}$
with an arbitrary fixed integer $r\geqslant 1$, provided that the three coefficients are all bounded. Note that the above bound is non-trivial as long as
$K>p^{\frac{1}{4}+\eta},\qquad KM>p^{\frac{1}{2}+\eta},\qquad N>p^{\eta}$
with some constant $\eta>0$.
Our work on $\mathfrak{T}(\bm{\alpha},\bm{\beta})$ is largely inspired by the recent work of Fouvry and Shparlinski [FS23] on quadrilinear character sums such as
$\mathfrak{Q}_{1}(\bm{\alpha},\bm{\beta})=\mathop{\sum\sum\sum\sum}_{1\leqslant r,s,u,v\leqslant x}\alpha_{r}\beta_{u}\chi(rs-uv),$
$\mathfrak{Q}_{2}(\bm{\gamma})=\mathop{\sum\sum\sum\sum}_{1\leqslant r,s,u,v\leqslant x}\gamma_{r,s,u}\chi(rs-uv),$
with bounded complex weights $\bm{\alpha}=(\alpha_{r})$, $\bm{\beta}=(\beta_{u})$ and $\bm{\gamma}=(\gamma_{r,s,u})$. It is shown in [FS23], that for any fixed $\eta>0$ and $x\geqslant p^{\frac{1}{8}+\eta}$, we have
$\mathfrak{Q}_{1}(\bm{\alpha},\bm{\beta})\ll x^{4-\kappa}\qquad\mbox{and}\qquad\mathfrak{Q}_{2}(\bm{\gamma})\ll x^{4}\frac{\log\log p}{\log p}$
where $\kappa>0$ and the implied constants depend only on $\eta$ (we refer to Section 1.4 for the exact definition of the symbol $\ll$’ and other standard notations).
1.4. Notation and conventions
We adopt the Landau symbol $A=O(B)$, and the Vinogradov symbol $A\ll B$, to mean $|A|\leqslant cB$ for some constant $c>0$. We also write $a\sim A$ to indicate that $A<a\leqslant 2A$ and $A\asymp B$ to indicate $A\ll B\ll A$.
For each complex weight $\bm{\alpha}=(\alpha_{m})_{m\in{\mathcal{M}}}$ and $\rho\geqslant 1$, we denote
(1.7) $\|\bm{\alpha}\|_{\rho}=\left(\sum_{m\in{\mathcal{M}}}|\alpha_{m}|^{\rho}\right)^{1/\rho}\qquad\mbox{and}\qquad\|\bm{\alpha}\|_{\infty}=\max_{m\in{\mathcal{M}}}|\alpha_{m}|.$
For a finite set ${\mathcal{S}}$ we use #$\#{\mathcal{S}}$ to denote its cardinality. The letter $p$, with or without subscripts, always denotes a prime number.
We denote by ${\mathbb{F}}_{p}$ the finite field of $p$ elements, which we identify by $\{0,1,\ldots,p-1\}$, and in what follows, we mix the usage of ${\mathbb{F}}_{p}$ and $\{0,1,\ldots,p-1\}$. We freely alternate between the language of finite fields and the language of congruences.
We also use ${\mathbb{N}}$ to denote that set of positive integers.
As usual,
(1.8) $\tau(k)=\#\{d\in{\mathbb{N}}:\leavevmode\nobreak\ d\mid k\}$
denotes the divisor function, and we repeatedly use the classical bound $\tau(k)=k^{o(1)}$ as $k\to+\infty$ (see, for example, in [IK04, Equation (1.81)]). For an integer $a$ coprime to $m$, $\overline{a}$ denotes the multiplicative inverse of $a\bmod m$, that is, $a\overline{a}\equiv 1\bmod m$, which should not be confused with the complex conjugate.
Given a function $f\in L^{1}({\mathbb{R}})$, that is, with a bounded $L^{1}$-norm over ${\mathbb{R}}$, the Fourier transform is defined by
$\widehat{f}(y)=\int_{\mathbb{R}}f(x)\mathrm{e}(-yx)\mathrm{d}x$
with $\mathrm{e}(z)=\exp(2\pi iz)$.
2. Main results
2.1. Bounds of multilinear character sums
Put
(2.1) $\mathscr{L}_{1}=(|a|K+|b|MN)\qquad\mbox{and}\qquad\mathscr{L}_{2}=(|a|KL+|b|MN).$
We prove two estimates for the sums $\mathfrak{T}(\bm{\alpha},\bm{\beta})$ given by (1.1), according to whether the weights $\alpha_{m}$ are identically $1$, which we write as $\bm{\alpha}\equiv\mathbf{1}$, or arbitrary.
Theorem 2.1.
Let $K,M,N>1$ and $p>\max\{K,M,N\}$ a large prime. Uniformly over the weights $\bm{\alpha}\equiv\mathbf{1},$ $\bm{\beta}=(\beta_{k,n})$ and integers $a,b$ with $\gcd(ab,p)=1$, we have
$|\mathfrak{T}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\beta}\|_{\infty}KMN\Delta_{1}\,p^{o(1)}$
for each positive integer $r$, where
$\Delta_{1}=(1+NK^{-1})^{\frac{1}{2r}}(1+\mathscr{L}_{1}MNp^{-1-\frac{1}{2r}})^{\frac{1}{2r}}\left(\frac{p^{1+\frac{1}{r}}}{M^{4}N^{4}}\right)^{\frac{1}{4r}}$
with $\mathscr{L}_{1}$ as defined by (2.1).
Remark 2.2.
Taking $N=K=1$ in Theorem 2.1, we recover the celebrated Burgess bound for short linear character sums, which shows oscillations of non-trivial multiplicative characters modulo $p$, in any interval with length at least $p^{\frac{1}{4}+\varepsilon}$ for any $\varepsilon>0$ and sufficiently large $p$. This is also known as the Burgess threshold (see [IK04, Theorem 12.26]).
Remark 2.3.
It is not easy to give a full description on the range of $(K,M,N)$ which is equivalent to $\Delta_{1}<1$. However, when $a=b=1$, we note that Theorem 2.1 is non-trivial if either of the following conditions
  • $p^{\frac{1}{4}+\eta}<MN\leqslant p^{\frac{1}{2}}$ and $N\leqslant K\leqslant p/(MN)$,
  • $MN\leqslant p^{\frac{1}{2}+\eta}\leqslant M^{2}NK$ and $N\geqslant K$,
holds with some fixed $\eta>0$. For example, we can take
(2.2) $K,M,N\sim p^{\frac{1}{8}+\eta},$
for small $\eta>0$. The lower bound for the values of $K,M,N$ given by (2.2), is the square root of the Burgess threshold.
We now use Theorem 2.1 to get a new bound, with a power saving, on the sums $\mathfrak{Q}_{2}(\bm{\gamma})$ from Section 1.3. Indeed, applying Theorem 2.1 with
$(m,n,k;a,b)\leftarrow(v,u,k;1,-1),\qquad(K,M,N)\leftarrow(x^{2},x,x)$
and
$\beta_{n,k}=\mathop{\sum\sum\sum}_{\begin{subarray}{c}1\leqslant r,s,u\leqslant x\\ rs=k,\,u=n\end{subarray}}\gamma_{r,s,u},$
we find the following.
Corollary 2.4.
Let $p$ be a large prime with $p>x^{2}\geqslant 1$. Uniformly over the weights $\bm{\gamma}=(\gamma_{r,s,u})$, we have
$|\mathfrak{Q}_{2}(\bm{\gamma})|\leqslant\|\bm{\gamma}\|_{\infty}x^{4-\frac{2}{r}}(1+x^{4}p^{-1-\frac{1}{2r}})^{\frac{1}{2r}}p^{\frac{1}{4r}+\frac{1}{4r^{2}}+o(1)}$
for each positive integer $r$.
We see that Corollary 2.4 gives a non-trivial bound for $\mathfrak{Q}_{2}(\bm{\gamma})$ with a power-saving as long as $x\geqslant p^{\frac{1}{8}+\eta}$ with some fixed $\eta>0$. And this also recovers the above bound for $\mathfrak{Q}_{1}(\bm{\alpha},\bm{\beta})$ with more general weights.
For arbitrary weights $\bm{\alpha}$ and $\bm{\beta}$, we have the following alternative bound.
Theorem 2.5.
Let $K,M,N>1$ and $p>\max\{K,M,N\}$ a large prime. Uniformly over the weights $\bm{\alpha}=(\alpha_{m})$, $\bm{\beta}=(\beta_{k,n})$ and integers $a,b$ with $\gcd(ab,p)=1$, we have
$|\mathfrak{T}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|_{2}M(NK)^{\frac{1}{2}}\Delta_{2}\,p^{o(1)}$
for each positive integer $r$, where
$\Delta_{2}=(1+KM^{-2})^{\frac{1}{4r}}(1+\mathscr{L}_{1}MNp^{-1-\frac{1}{r}})^{\frac{1}{2r}}\left(\frac{p^{1+\frac{1}{r}}}{K(MN)^{2}}\right)^{\frac{1}{4r}}+M^{-\frac{1}{2}}$
with $\mathscr{L}_{1}$ as defined by (2.1).
Remark 2.6.
Fix $a=b=1$, then Theorem 2.5 is non-trivial as long as
$M>p^{\eta},\qquad K(MN)^{2}>(1+KM^{-2})p^{1+\eta},\qquad MN(K+MN)<p,$
holds with some fixed $\eta>0$. For example, we can take
$K,M,N\sim p^{\frac{1}{5}+\eta},$
which has to be compared with (2.2) and also with the Burgess threshold.
As an application of the above bounds for $\mathfrak{T}(\bm{\alpha},\bm{\beta})$, one is allowed to address a modulo $p$ version of a question of Iwaniec and Sárközy [IS87] about distances between product sets and squares. We present such an application in § 2.2.
One can derive some bounds on such sums from Theorems 2.1 and 2.5 with a trivial summation over $a$, which plays the role of $\ell$ in $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$. However we may obtain a more precise bound as follows.
Theorem 2.7.
Let $K,L,M,N>1$ and $p>\max\{K,L,M,N\}$ a large prime. Uniformly over the weights $\bm{\alpha}=(\alpha_{\ell,m})$ and $\bm{\beta}=(\beta_{k,n})$, we have
$|\mathfrak{Q}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|_{2}LM(KN)^{\frac{1}{2}}p^{o(1)}\cdot\Delta_{3},$
for each positive integer $r$, where
$\Delta_{3}=(KLMN)^{-\frac{3}{4r}}$ $\left(\frac{M}{K}+1\right)^{\frac{1}{2r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{p^{1+\frac{1}{r}}}\right)^{\frac{1}{2r}}\mathscr{L}_{2}^{\frac{1}{2r}}p^{\frac{1}{4r}+\frac{1}{2r^{2}}}$
$\qquad+(KN)^{-\frac{1}{2}}p^{\frac{1}{2r}}+(MN)^{-\frac{1}{2}}+(LM)^{-\frac{1}{2}}+p^{-\frac{1}{2}}$
with $\mathscr{L}_{2}$ as defined by (2.1).
We now analyse when Theorem 2.7 wins over the trivial bound
$|\mathfrak{Q}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|_{2}(LM)(KN)^{\frac{1}{2}}.$
Corollary 2.8.
Fix $a=b=1$. For any $\eta>0$ there is some $\kappa>0$ such that if
(2.3) $\begin{cases}KL\ll MN\leqslant KLp^{\frac{1}{3}-\eta},\\ K^{5}L^{3}N\geqslant Mp^{1+\eta},\\ (M/K)^{5}(N/L)^{3}\leqslant p^{1-\eta},\\ (KL)^{3}MN\geqslant p^{1+\eta},\\ KN\geqslant p^{\eta},\\ MN\geqslant p^{\eta},\\ LM\geqslant p^{\eta},\end{cases}$
then
$|\mathfrak{Q}(\bm{\alpha},\bm{\beta})|\leqslant\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|LM(KN)^{\frac{1}{2}}p^{-\kappa}.$
To see this, we write the quantity $\Delta_{3}$ in an obvious manner as
(2.4) $\Delta_{3}=\widetilde{\Delta}_{3}+(KN)^{-\frac{1}{2}}p^{\frac{1}{2r}}+(MN)^{-\frac{1}{2}}+(LM)^{-\frac{1}{2}}+p^{-\frac{1}{2}}$
corresponding to five terms in its definition. By taking $r$ sufficiently large, each of the last four terms in (2.4) is at most $p^{-\kappa}$ for some $\kappa>0$, provided that the last three conditions in (2.3) hold. It remains to check that $\widetilde{\Delta}_{3}\ll p^{-\kappa}$. We now assume $KL\ll MN$ so that $\mathscr{L}_{2}=KL+MN\ll MN$. First, rising $\widetilde{\Delta}_{3}$ to the power $2r$ and expanding it, we obtain four terms. Each of these terms is at most $p^{-\kappa}$ as long as $KL\ll MN$ and
$\max\left\{(MN)^{2}p^{-\frac{1}{2}},\,\dfrac{M}{K}p^{\frac{1}{2}+\frac{1}{r}},\,\dfrac{M^{3}N^{2}}{Kp^{\frac{1}{2}}},\,p^{\frac{1}{2}+\frac{1}{r}}\right\}\ll(KL)^{\frac{3}{2}}(MN)^{\frac{1}{2}}p^{-2r\kappa}.$
These leads to the remaining inequalities in (2.3), after choosing $r$ to be large enough.
In particular, we see from Corollary 2.8 that we have a non-trivial bound on $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$ provided that
$K=L=M=N>p^{\frac{1}{8}+\eta}.$
Observe that one more time the exponent $1/8$ appears as a threshold, see Remark 2.3, however this does not seem to follow from Theorem 2.1.
2.2. Consequences and applications
Theorems 2.1 and 2.5 describe general situations since the sequence $\bm{\beta}$ satisfies no particular hypotheses. For instance, if we choose $\bm{\alpha}$ and $\bm{\beta}$ in the landscape of multiplicative or additive characters modulo $p$, we obtain, as a direct consequence of Theorem 2.5, the following bound for triple character sums on short initial segments. To proceed, we use $\psi$ to denote a non-trivial additive character of ${\mathbb{F}}_{p}$.
Corollary 2.9.
Fix $P$ a non-zero rational function in ${\mathbb{F}}_{p}(X)$ and $Q$ a non-zero rational function in ${\mathbb{F}}_{p}(X,Y)$. For every fixed $\eta>0$ there exists some $\kappa>0$, such that
$\sum_{k\leqslant K}\sum_{m\leqslant M}\sum_{n\leqslant N}\chi(k+mn)\psi(P(m)+Q(k,n))\ll KMNp^{-\kappa}$
and
$\sum_{k\leqslant K}\sum_{m\leqslant M}\sum_{n\leqslant N}\chi((k+mn)P(m)Q(k,n))\ll KMNp^{-\kappa}$
hold uniformly for
$p^{\frac{1}{5}+\eta}\leqslant K,\,M,\,N<p^{\frac{1}{4}-\eta}.$
Remark 2.10.
Theorem 2.7 easily leads to an obvious quadrilinear version of Corollary 2.9, but with $K,L,M$ and $N$ of size at least $p^{1/8+\eta}$ for any $\eta>0$.
Furthermore, Iwaniec and Sárközy [IS87] have considered the following multiplicative hybrid problem with positive integers: given two arbitrary subsets ${\mathcal{M}}$ and ${\mathcal{N}}$, how close is the product $mn$ to a square in ${\mathbb{Z}}$ with $(m,n)\in{\mathcal{M}}\times{\mathcal{N}}$? A special case of their general result asserts that for any ${\mathcal{M}},{\mathcal{N}}\subseteq[N,2N]$ with #,#$\#{\mathcal{M}},\#{\mathcal{N}}\gg N$, there exist $(m,n,\ell)\in{\mathcal{M}}\times{\mathcal{N}}\times{\mathbb{N}}$ satisfying
$\frac{mn-\ell^{2}}{\ell}\ll(N/\log N)^{-\frac{1}{2}}.$
Iwaniec and Sárközy [IS87] have also conjectured that the upper bound might be replaced by $N^{-1+o(1)}$.
Theorem 2.5 allows us to study a modular analogue of the above result of Iwaniec and Sárközy, for which the above conjectural bound $N^{-1+o(1)}$ can be realised in some particular cases. To set up, we consider subsets ${\mathcal{M}},{\mathcal{N}}\subseteq[1,p]$, we examine the distance between $mn$, with $(m,n)\in{\mathcal{M}}\times{\mathcal{N}}$, and quadratic residues modulo $p$.
Theorem 2.11.
Fix a positive integer $r$, two real numbers $10<c_{0}<$ 1 and $\eta>0$. Then there exists a constant $P$, depending only on $(r,c_{0},\eta)$, such that
  • for every prime $p\geqslant P$,
  • for every $M$ and $N$ satisfying
    (2.5) $M^{4}N^{2}\geqslant p^{1+\frac{1}{r}+\eta},$
  • for every subsets ${\mathcal{M}},{\mathcal{N}}\subseteq[1,p]$ with
    (2.6) ${\mathcal{M}}\subseteq[M,2M],\quad{\mathcal{N}}\subseteq[N,2N],\quad\#{\mathcal{M}}\geqslant c_{0}M,\quad\#{\mathcal{N}}\geqslant c_{0}N,$
there exist $m\in{\mathcal{M}}$, $n\in{\mathcal{N}}$, and some positive integer $k$ satisfying
$k\leqslant p^{1+\frac{1}{r}+\eta}(MN)^{-2}+p^{\eta}$
such that $mn+k$ is a quadratic residue modulo $p$.
Actually, the proof which is given in § 7 produces a lower bound for the cardinality of the set of triples we are interested in
#$\#\{(m,n,k):\leavevmode\nobreak\ m\in{\mathcal{M}},\,n$ $\in{\mathcal{N}},\,k\leqslant K\text{ and }mn+k$
[…]
where $K\sim p^{1+\frac{1}{r}+\eta}(MN)^{-2}$.
To illustrate Theorem 2.11, choose $M=N\sim p^{\frac{1}{5}+\delta}$ (with $\delta>0$ very small) and ${\mathcal{M}}$ and ${\mathcal{N}}$ subsets of $[N,2N]$ such that #,#$\#{\mathcal{M}},\#{\mathcal{N}}\geqslant\delta N$. Then, for sufficiently large $p$, there exists a positive integer $k\leqslant p^{\frac{1}{5}}$ and $(m,n)\in{\mathcal{M}}\times{\mathcal{N}}$ such that $mn+k$ is a quadratic residue modulo $p$. Recall that there must exist a square of an integer between $t$ and $t+O(\sqrt{t})$ for all $t>1$ (namely, $\lceil\sqrt{t}\rceil^{2}$ is such a square) and that essentially, nothing better than the upper bound $p^{\frac{1}{4}+o(1)}$ is proved for the distance between two consecutive quadratic residues modulo $p$. The above illustration shows that one can do much better if squares are replaced by quadratic residues modulo $p$.
The proof of Theorem 2.11 can be easily generalized in many aspects. For instance one can relax the hypothesis (2.6) about the cardinalities of ${\mathcal{M}}$ and ${\mathcal{N}}$ by choosing
(2.7) $\begin{cases}\alpha_{m}={\bf 1}_{[M,2M]\cap{\mathcal{P}}}(m),\\ \beta_{k,n}={\bf 1}_{[K,2K]\cap{\mathcal{P}}}(k)\cdot{\bf 1}_{[N,2N]\cap{\mathcal{P}}}(n),\end{cases}$
where ${\mathcal{P}}$ denotes the set of all primes, and ${\bf 1}_{S}$ is the indicator function of a set $S$. One then deduces from Theorem 2.5 the following result by counting primes using Chebyshev’s inequalities.
Corollary 2.12.
Fix $\eta>0$. For every sufficiently large $p$, there exist three primes $p_{1},p_{2}$ and $p_{3}$ such that
  • $p_{1},p_{2},p_{3}\sim p^{{\frac{1}{5}}+\eta}$,
  • $p_{1}p_{2}+p_{3}$ is a quadratic non-residue modulo $p$.
As we have mentioned, Corollary 2.12 follows directly from Theorem 2.5 and it can be modified in various ways, for example one can impose different arithmetic restrictions on the primes $p_{1},p_{2},p_{3}$. For example, one further imposes that the three shifted primes $p_{1}+2,p_{2}+2,p_{3}+2$ have at most two prime factors. This statement is deduced from the famous work of Chen [Ch73] about the twin prime conjecture that we write under the form of the following inequality
#$\#\{p\leqslant x:\leavevmode\nobreak\ p+2\text{ has at most two prime factors}\}\gg x(\log 2x)^{-2},$
for $x\geqslant 2$. One can also appeal to [Fo82, Corollaire 2] to introduce a more involved definition of the coefficient $\beta_{k,n}$ (see (2.7)) leading to a new version of Corollary 2.12, where we impose on $p_{1}p_{3}+2$ to have at most two prime factors.
Similarly, Theorem 2.7 leads to the following result (which can also be modified along the lines mentioned in the above).
Corollary 2.13.
Fix $\eta>0$. For every sufficiently large $p$, there exist four primes $p_{1},p_{2},p_{3}$ and $p_{4}$ such that
  • $p_{1},p_{2},p_{3},p_{4}\sim p^{{\frac{1}{8}}+\eta}$,
  • $p_{1}p_{2}+p_{3}p_{4}$ is a quadratic non-residue modulo $p$.
Another application of Theorem 2.7 concerns sums with the divisor function (see (1.8) below)
$S(U,V)=\sum_{u\leqslant U}\sum_{v\leqslant V}\tau(u)\tau(v)\chi(u-v).$
These sums are two-dimensional analogues of the sum
$S_{a}(U)=\sum_{u\leqslant U}\tau(u)\chi(u+a),$
which has been studied in a number of works [BGKS13, Ch10, Ka71, Ka00]. In particular, it is shown in [BGKS13, Theorem 27] that for any $\eta>0$ there is some $\kappa>0$ such that for $U>p^{\frac{1}{3}+\eta}$, uniformly over integers $a$ with $\gcd(a,p)=1$, one has
$S_{a}(U)\ll Up^{-\kappa}.$
It is easy to see that Theorem 2.7 combined with the standard completing technique (see [IK04, Section 12.2]) implies the following result.
Corollary 2.14.
For any fixed $\eta>0$, there is some $\kappa>0$ such that for $U,V>p^{\frac{1}{4}+\eta}$ we have
$S(U,V)\ll UVp^{-\kappa}.$
One can also use Theorem 2.7 to estimate a rich variety of other quantities, for example, sums over primes
$W(x)=\sum_{p_{1},p_{2},p_{3},p_{4}\leqslant x}\left(\frac{p_{1}}{p_{3}}\right)\left(\frac{p_{2}}{p_{4}}\right)\chi\left(p_{1}p_{2}-p_{3}p_{4}\right)$
with weights given by Legendre symbols. Theorem 2.7 implies a bound on $W(x)$ with a power saving, provided $x\geqslant p^{\frac{1}{8}+\eta}$ for any fixed $\eta>0$.We also note a variety of bounds on characters over various arithmetic sequences can be found in a very informative survey of Karatsuba [Ka08].
In closing this section, we would like to state a striking application of our general bounds in Theorem 2.7 to character sums involving Farey fractions. For $R\geqslant 2$ let
[…]
be the set of Farey fractions of order $R$. This can be embedded in ${\mathbb{F}}_{p}$ in a canonical way $r/s\mapsto r\overline{s}\pmod{p}$, where $\overline{s}$ is the multiplicative inverse of $s$ modulo $p$ (which is well-defined for $R<p$ and in fact is injective for $R<p^{\frac{1}{2}}$).
By virtue of the multiplicativity of $\chi$, in the form
$\alpha_{\ell,m}\beta_{k,n}\chi(\ell\overline{m}-\overline{k}n)=\left(\alpha_{\ell,m}\overline{\chi}(m)\right)\left(\beta_{k,n}\overline{\chi}(k)\right)\chi(k\ell-mn)$
we may derive the following consequence directly from Theorem 2.7.
Corollary 2.15.
Let $R\geqslant 2$ and $\xi_{\rho},\zeta_{\varrho}$ be bounded weights supported on ${\mathcal{F}}(R)$. Then for any fixed $\eta>0$, there is some $\kappa>0$ such that for $p^{\frac{1}{8}+\eta}\leqslant R<p^{\frac{1}{2}}$ we have
$\mathop{\sum\sum}_{\rho,\varrho\in{\mathcal{F}}(R)}\xi_{\rho}\zeta_{\varrho}\chi(\rho-\varrho)\ll R^{4}p^{-\kappa}.$
Note that the use of bilinear bounds (1.5) and (1.6) would lead to a much more restrictive condition $p^{\frac{1}{4}+\eta}\leqslant R<p^{\frac{1}{2}}$ in Corollary 2.15.
3. Preliminaries
3.1. Moments of some character sums
The following result is a consequence of Riemann Hypothesis for curves over finite fields due to Weil [We48]. It is a slight extension of a classical bound due to Davenport and Erdős [DE52], which corresponds to the case when ${\mathcal{A}}$ is an interval and $\bm{\gamma}=1$. The proof of Lemma 3.1 below is identical, and is also a part of the arguments in Appendix A.
Lemma 3.1.
Let ${\mathcal{A}}\subseteq{\mathbb{F}}_{p}$ be a subset of cardinality $A$, and $\chi$ a non-trivial multiplicative character of ${\mathbb{F}}_{p}^{\times}$. For each positive integer $r$ and any complex-valued weight $\bm{\gamma}=(\gamma_{a})$ with $\|\bm{\gamma}\|_{\infty}\leqslant 1$, we have
$\sum_{x\in{\mathbb{F}}_{p}}\left|\sum_{a\in{\mathcal{A}}}\gamma_{a}\chi(x+a)\right|^{2r}\leqslant(2r)^{r}(A^{r}p+A^{2r}p^{\frac{1}{2}}).$
We also need the following version of Lemma 3.1, which again essentially repeats the argument in [DE52] but uses Weil’s bound for multiplicative character sums with polynomials; see, for example, [IK04, Theorem 11.23].
Lemma 3.2.
Let ${\mathcal{A}}\subseteq{\mathbb{F}}_{p}$ be a subset of cardinality $A$, and $\chi_{1},\chi_{2}$ two non-trivial multiplicative characters of ${\mathbb{F}}_{p}^{\times}$. For each positive integer $r$ and any complex-valued weight $\bm{\gamma}=(\gamma_{a})$ with $\|\bm{\gamma}\|_{\infty}\leqslant 1$, we have
$\mathop{\sum\sum}_{x,y\in{\mathbb{F}}_{p}}\left|\sum_{a\in{\mathcal{A}}}\gamma_{a}\chi_{1}(x+a)\chi_{2}(y+a)\right|^{2r}\leqslant(2r)^{r}(A^{r}p^{2}+2rA^{2r}p).$
Proof.
Denote by $S$ the quantity in question. Opening the power, we write
$S\leqslant\mathop{\sum\sum}_{\mathbf{a},\mathbf{b}\in{\mathcal{A}}^{r}}|S(\mathbf{a},\mathbf{b};\chi_{1})S(\mathbf{a},\mathbf{b};\chi_{2})|,$
where for $\mathbf{a}=(a_{1},\cdots,a_{r})\in{\mathcal{A}}^{r}$, $\mathbf{b}=(b_{1},\cdots,b_{r})\in{\mathcal{A}}^{r}$ and each non-trivial multiplicative character $\chi$ of ${\mathbb{F}}_{p}^{\times}$,
$S(\mathbf{a},\mathbf{b};\chi)=\sum_{x\in{\mathbb{F}}_{p}}\prod_{1\leqslant j\leqslant r}\chi(x+a_{j})\overline{\chi(x+b_{j})}.$
The subsequent treatment is uniform in all non-trivial multiplicative characters $\chi$ of ${\mathbb{F}}_{p}^{\times}$. If the coordinates of $\mathbf{a},\mathbf{b}$ appear in pairs (with necessary permutations), we appeal to the trivial bound
$|S(\mathbf{a},\mathbf{b};\chi)|\leqslant p.$
Note that the number of such tuples $(\mathbf{a},\mathbf{b})$ is at most $r\binom{2r}{r}A^{r}\leqslant(2rA)^{r}$. For the remaining $\mathbf{a},\mathbf{b}$, we can apply Weil’s bound for complete character sums (see [IK04, Corollary 11.24]), getting
$|S(\mathbf{a},\mathbf{b};\chi)|\leqslant 2rp^{\frac{1}{2}}.$
[…]
3.2. Bounds of some GCD sums
We need the following estimate.
Lemma 3.3.
Let $a,b$ be non-zero integers. Let $A,B,K,L,M,N,U,W\geqslant 1$ with $A\ll|a|KL+|b|MN$ and $B\ll LU+MW$. Then we have
$\mathop{\sum_{k\leqslant K}\mathop{\sum\sum}_{\ell_{1},\ell_{2}\leqslant L}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\sum_{n\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\begin{subarray}{c}ak\ell_{\nu}+bm_{\nu}n=a_{\nu},\leavevmode\nobreak\ a\ell_{\nu}u+bm_{\nu}w=b_{\nu},\ \nu=1,2\\ |a_{1}|\sim A,\leavevmode\nobreak\ |b_{2}|\sim B,\leavevmode\nobreak\ kw\neq nu\end{subarray}}\gcd(a_{1},b_{1})\gcd(a_{2},b_{2})\leqslant ABLM(KW+NU)\Upsilon^{o(1)}$
with $\Upsilon=|ab|ABKLMNUW$.
Proof.
Denote by $G$ the above quantity in question. Writing $g_{1}=\gcd(a_{1},b_{1})$ and $g_{2}=\gcd(a_{2},b_{2})$ we have the inequality
$G$ $\leqslant\mathop{\sum\sum}_{g_{1}\ll A,\,g_{2}\ll B}g_{1}g_{2}\mathop{\sum_{k\leqslant K}\mathop{\sum\sum}_{\ell_{1},\ell_{2}\leqslant L}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\sum_{n\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\begin{subarray}{c}ak\ell_{\nu}+bm_{\nu}n\equiv a\ell_{\nu}u+bm_{\nu}w\equiv 0\bmod{g_{\nu}},\ \nu=1,2\\ |ak\ell_{1}+bm_{1}n|\sim A,\leavevmode\nobreak\ |a\ell_{2}u+bm_{2}w|\sim B,\leavevmode\nobreak\ kw\neq nu\end{subarray}}1.$
The congruences in the summation imply
$g_{\nu}\mid m_{\nu}w(ak\ell_{\nu}+bm_{\nu}n)-m_{\nu}n(a\ell_{\nu}u+bm_{\nu}w)=a\ell_{\nu}m_{\nu}(kw-nu)$
for $\nu=1,2$, so that we can decompose $g_{\nu}$ (in a not necessarily unique way) as
$g_{\nu}=d_{\nu}e_{\nu}f_{\nu}$
where
$d_{\nu}\mid a\ell_{\nu},\qquad e_{\nu}\mid m_{\nu},\qquad f_{\nu}\mid(kw-nu).$
In particular, we have
$\operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU.$
Therefore, we have
$G\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}e_{1}f_{1}\ll A\\ d_{2}e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}\mathop{\sum_{k\leqslant K}\sum_{n\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\operatorname{lcm}[f_{1},f_{2}]\mid(kw-nu)\neq 0}\mathop{\mathop{\sum\sum\sum\sum}_{\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M}}_{\begin{subarray}{c}ak\ell_{1}+bm_{1}n\equiv 0\bmod{d_{1}e_{1}f_{1}}\\ a\ell_{2}u+bm_{2}w\equiv 0\bmod{d_{2}e_{2}f_{2}}\\ \ell_{\nu}\equiv 0\bmod{d_{\nu}},\,m_{\nu}\equiv 0\bmod{e_{\nu}},\ \nu=1,2\\ |ak\ell_{1}+bm_{1}n|\sim A,\leavevmode\nobreak\ |a\ell_{2}u+bm_{2}w|\sim B\end{subarray}}1.$
Making the change of variables $\ell_{\nu}\to d_{\nu}\ell_{\nu}$, $m_{\nu}\to e_{\nu}m_{\nu}$, $\nu=1,2$, we obtain
$G\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}e_{1}f_{1}\ll A\\ d_{2}e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}\mathop{\sum_{k\leqslant K}\sum_{k\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\operatorname{lcm}[f_{1},f_{2}]\mid(kw-nu)\neq 0}\mathop{\mathop{\sum\sum\sum\sum}_{d_{1}\ell_{1},\,d_{2}\ell_{2}\leqslant L,\leavevmode\nobreak\ e_{1}m_{1},\,e_{2}m_{2}\leqslant M}}_{\begin{subarray}{c}ad_{1}\ell_{1}k+be_{1}m_{1}n\equiv 0\bmod{d_{1}e_{1}f_{1}}\\ ad_{2}\ell_{2}u+be_{2}m_{2}w\equiv 0\bmod{d_{2}e_{2}f_{2}}\\ |ad_{1}\ell_{1}k+be_{1}m_{1}n|\sim A,\leavevmode\nobreak\ |ad_{2}\ell_{2}u+be_{2}m_{2}w|\sim B\end{subarray}}1.$
Note that the congruences
$ad_{1}\ell_{1}k+be_{1}m_{1}n\equiv 0\bmod{d_{1}e_{1}f_{1}}\qquad\mbox{and}\qquad ad_{2}\ell_{2}u+be_{2}m_{2}w\equiv 0\bmod{d_{2}e_{2}f_{2}}$
imply
$e_{1}\mid ad_{1}\ell_{1}k,\qquad d_{2}\mid be_{2}m_{2}w,$
and
$bm_{1}n\equiv-ad_{1}\ell_{1}k/e_{1}\bmod{d_{1}f_{1}},\qquad a\ell_{2}u\equiv-be_{2}m_{2}w/d_{2}\bmod{e_{2}f_{2}},$
respectively. It then follows that
$G$ $\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}e_{1}f_{1}\ll A\\ d_{2}e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}\mathop{\sum_{k\leqslant K}\sum_{n\leqslant N}\sum_{u\sim U}\sum_{w\sim W}}_{\operatorname{lcm}[f_{1},f_{2}]\mid(kw-nu)\neq 0}\mathop{\mathop{\sum\sum\sum\sum}_{d_{1}\ell_{1},d_{2}\ell_{2}\leqslant L,\leavevmode\nobreak\ e_{1}m_{1},e_{2}m_{2}\leqslant M}}_{\begin{subarray}{c}e_{1}\mid ad_{1}\ell_{1}k,\leavevmode\nobreak\ bm_{1}n\equiv-ad_{1}\ell_{1}k/e_{1}\bmod{d_{1}f_{1}}\\ d_{2}\mid be_{2}m_{2}w,\leavevmode\nobreak\ a\ell_{2}u\equiv-be_{2}m_{2}w/d_{2}\bmod{e_{2}f_{2}}\\ |ad_{1}\ell_{1}k+be_{1}m_{1}n|\sim A,\leavevmode\nobreak\ |ad_{2}\ell_{2}u+be_{2}m_{2}w|\sim B\end{subarray}}1$
$\leqslant\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}d_{1}e_{2}f_{1}f_{2}\sup_{\delta\in{\mathbb{Z}}}\mathop{\sum_{k\leqslant K}\sum_{w\sim W}}_{kw\equiv\delta\bmod{\operatorname{lcm}[f_{1},f_{2}]}}\mathop{\sum\sum}_{\begin{subarray}{c}d_{2}\ll B/(e_{2}f_{2})\\ e_{1}\ll A/(d_{1}f_{1})\end{subarray}}d_{2}e_{1}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\quad\times\sum_{n\leqslant N}\sum_{u\sim U}\mathop{\mathop{\sum\sum\sum\sum}_{d_{1}\ell_{1},d_{2}\ell_{2}\leqslant L,\leavevmode\nobreak\ e_{1}m_{1},e_{2}m_{2}\leqslant M}}_{\begin{subarray}{c}e_{1}\mid ad_{1}\ell_{1}k,\leavevmode\nobreak\ bm_{1}n\equiv-ad_{1}\ell_{1}k/e_{1}\bmod{d_{1}f_{1}}\\ d_{2}\mid be_{2}m_{2}w,\leavevmode\nobreak\ a\ell_{2}u\equiv-be_{2}m_{2}w/d_{2}\bmod{e_{2}f_{2}}\\ |ad_{1}\ell_{1}k/e_{1}+bm_{1}n|\sim A/e_{1},\leavevmode\nobreak\ |a\ell_{2}u+be_{2}m_{2}w/d_{2}|\sim B/d_{2}\end{subarray}}1.$
To proceed, we group the variables $m_{1},n$, and $\ell_{2},u$, separately, so that we need to count the number of tuples $(r,s)$ with
$|ad_{1}\ell_{1}k/e_{1}+br|\sim A/e_{1},\leavevmode\nobreak\ |as+be_{2}m_{2}w/d_{2}|\sim B/d_{2}$
and
$br\equiv-d_{1}\ell_{1}k/e_{1}\bmod{d_{1}f_{1}},\qquad as\equiv-e_{2}m_{2}w/d_{2}\bmod{e_{2}f_{2}}.$
It is clear that the number $T$ of such tuples $(r,s)$ satisfies
$T\ll\left(1+\frac{A\gcd(b,d_{1}f_{1})}{bd_{1}e_{1}f_{1}}\right)\left(1+\frac{B\gcd(a,e_{2}f_{2})}{ad_{2}e_{2}f_{2}}\right)\ll\frac{AB}{d_{1}d_{2}e_{1}e_{2}f_{1}f_{2}}.$
This leads us to
$G\ll AB\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}\sup_{\delta\in{\mathbb{Z}}}\mathop{\sum_{k\leqslant K}\sum_{w\sim W}}_{kw\equiv\delta\bmod{\operatorname{lcm}[f_{1},f_{2}]}}\mathop{\sum\sum}_{\begin{subarray}{c}d_{2}\ll B/(e_{2}f_{2})\\ e_{1}\ll A/(d_{1}f_{1})\end{subarray}}\mathop{\sum\sum}_{\begin{subarray}{c}\ell_{1}\leqslant L/d_{1},\leavevmode\nobreak\ m_{2}\leqslant M/e_{2}\\ d_{2}\mid be_{2}m_{2}w,\leavevmode\nobreak\ e_{1}\mid ad_{1}\ell_{1}k\end{subarray}}1.$
In what follows, we would like to sum over $d_{2},e_{1}$ firstly, and then $\ell_{1},m_{2}$, so that
$G\leqslant ABLM\Upsilon^{o(1)}\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}\frac{1}{d_{1}e_{2}}\sup_{\delta\in{\mathbb{Z}}}\mathop{\sum_{k\leqslant K}\sum_{w\sim W}}_{kw\equiv\delta\bmod{\operatorname{lcm}[f_{1},f_{2}]}}1.$
Again by grouping variables, we arrive at
$G$ $\leqslant ABLM)\Upsilon^{o(1)}\mathop{\sum\sum}_{\begin{subarray}{c}d_{1}f_{1}\ll A\\ e_{2}f_{2}\ll B\\ \operatorname{lcm}[f_{1},f_{2}]\leqslant KW+NU\end{subarray}}\frac{1}{d_{1}e_{2}}\left(\frac{KW}{\operatorname{lcm}[f_{1},f_{2}]}+1\right)$
$\leqslant ABLM\Upsilon^{o(1)}\sum_{d_{1}\ll A}\frac{1}{d_{1}}\sum_{e_{1}\ll A}\frac{1}{e_{2}}\sum_{f\ll KW+NU}\left(\frac{KW}{f}+1\right)$
$\leqslant ABLM(KW+NU)\Upsilon^{o(1)},$
[…]
4. Proof of Theorem 2.1
4.1. Preparations
Throughout this section, we assume that $\bm{\alpha}$ is identically 1 on its support. Suppose without loss of generality that $\|\bm{\beta}\|_{\infty}\leqslant 1$ and $M$ is a positive integer. Following an approach of Friedlander and Iwaniec [FI93], we define the function $w\in L^{1}({\mathbb{R}})$ as
$w(x)=\begin{cases}\min\{x,\leavevmode\nobreak\ 1,\leavevmode\nobreak\ M+1-x\},&\text{for\leavevmode\nobreak\ }x\in[0,M+1],\\ 0,&\text{otherwise}.\end{cases}$
The integration by parts implies
(4.1) $\widehat{w}(\xi)\ll\min\{M,|\xi|^{-1},|\xi|^{-2}\}.$
4.2. Amplification
The above function $w$ allows us to run the summation over $m$ to the whole set ${\mathbb{Z}}$. That is,
$\mathfrak{T}(\bm{\alpha},\bm{\beta})=\sum_{k\leqslant K}\sum_{m\in{\mathbb{Z}}}\sum_{n\leqslant N}w(m)\beta_{k,n}\chi(ak+bmn).$
Furthermore, for all integers $u,v$ we have
$\mathfrak{T}(\bm{\alpha},\bm{\beta})=\sum_{k\leqslant K}\sum_{m\in{\mathbb{Z}}}\sum_{n\leqslant N}w(m+uv)\beta_{k,n}\chi(ak+b(m+uv)n).$
We choose two real positive parameters $U$ and $V$ with
(4.2) $UV=\frac{1}{4}M.$
By Fourier inversion and summing over integers $u$ and $v$ with $u\sim U$, $v\sim V$ we have the amplified expression
$\mathfrak{T}(\bm{\alpha},\bm{\beta})\ll\frac{1}{UV}\sum_{k\leqslant K}\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{u\sim U}\int_{\mathbb{R}}|\widehat{w}(\xi)|\left|\sum_{v\sim V}\mathrm{e}(uv\xi)\chi(\overline{un}(ak+bmn)+bv)\right|\mathrm{d}\xi.$
Making the change of variable $\xi\rightarrow\xi/u$ and replacing $UV$ with $M$, in view of (4.1), it follows that
(4.3) $\begin{split}\mathfrak{T}(\bm{\alpha},\bm{\beta})&\ll\frac{1}{M}\sum_{k\leqslant K}\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{u\sim U}\int_{\mathbb{R}}\left|\widehat{w}\left(\frac{\xi}{u}\right)\right|\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(\overline{un}(ak+bmn)+bv)\right|\frac{\mathrm{d}\xi}{u}\\ &\ll\frac{1}{M}\int_{\mathbb{R}}\min\left\{\frac{M}{U},\frac{1}{|\xi|},\frac{U}{|\xi|^{2}}\right\}\\ &\qquad\qquad\qquad\times\sum_{k\leqslant K}\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{u\sim U}\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(\overline{un}(ak+bmn)+bv)\right|\mathrm{d}\xi.\end{split}$
This yields
$\mathfrak{T}(\bm{\alpha},\bm{\beta})\ll\frac{\log M}{M}\sum_{k\leqslant K}\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{u\sim U}\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(\overline{un}(ak+bmn)+bv)\right|$
for some $\xi\in{\mathbb{R}}$ (for which the multiple sum in (4.3) achieves its largest possible value).
Note that we may further restrict our summation to triples $(k,m,n)$ with $p\nmid(ak+bmn)$ up to an error term at most
$p^{o(1)}\mathop{\sum_{k\leqslant K}\sum_{\ell\leqslant MN}}_{ak+b\ell\equiv 0\bmod p}1\leqslant\left(\frac{MN}{p}+1\right)Kp^{o(1)}.$
Therefore,
$\mathfrak{T}(\bm{\alpha},\bm{\beta})\ll\frac{\log M}{M}\mathop{\sum_{k\leqslant K}\sum_{m\leqslant M}\sum_{n\leqslant N}}_{p\nmid(ak+bmn)}\sum_{u\sim U}\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(\overline{un}(ak+bmn)+bv)\right|$
$+\left(\frac{MN}{p}+1\right)$ $Kp^{o(1)}.$
4.3. Regrouping, counting and Weil’s bound
Put
$\varrho(x)=\mathop{\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{k\leqslant K}\sum_{u\sim U}}_{ak+bmn\equiv unx\not\equiv 0\bmod p}1$
for $x\bmod p$. We now have
$\mathfrak{T}(\bm{\alpha},\bm{\beta})\ll\frac{\log M}{M}\sum_{x\bmod p}\varrho(x)\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(x+bv)\right|+\left(\frac{MN}{p}+1\right)Kp^{o(1)}.$
Applying the Hölder inequality with $r\geqslant 1$, we obtain
(4.4) $\mathfrak{T}(\bm{\alpha},\bm{\beta})\ll\frac{\log M}{M}{\mathcal{I}}_{1}^{1-\frac{1}{r}}({\mathcal{I}}_{2}\cdot{\mathcal{I}}_{3})^{\frac{1}{2r}}+\left(\frac{MN}{p}+1\right)Kp^{o(1)}$
with
${\mathcal{I}}_{j}=\sum_{x\bmod p}\varrho(x)^{j},\qquad j=1,2,$
and
${\mathcal{I}}_{3}=\sum_{x\bmod p}\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(x+bv)\right|^{2r}.$
Trivially we have
(4.5) ${\mathcal{I}}_{1}\ll KMNU.$
Regarding ${\mathcal{I}}_{2}$, we refer to the following lemma.
Lemma 4.1.
With the above notation, we have
(4.6) ${\mathcal{I}}_{2}\leqslant KMU(K+N)\left(1+\frac{\left(|a|K+|b|MN\right)UN}{p}\right)p^{o(1)}.$
Proof.
It is easy to see that ${\mathcal{I}}_{2}$ is the number of 8–tuples
$(k,\widetilde{k},m,\widetilde{m},n,\widetilde{n},u,\widetilde{u})\in{\mathbb{N}}^{8}$
satisfying the conditions
$k,\widetilde{k}\leqslant K,\quad m,$ $\,\widetilde{m}\leqslant M,\quad n,\widetilde{n}\leqslant N,\quad u,\widetilde{u}\sim U,$
$(ak+bmn)\widetilde{n}\widetilde{u}$ $\equiv(a\widetilde{k}+b\widetilde{m}\widetilde{n})nu\not\equiv 0\pmod{p},$
which is bounded by the number of 9–tuples
$(k,\widetilde{k},m,\widetilde{m},n,\widetilde{n},u,\widetilde{u},t)\in{\mathbb{N}}^{8}\times{\mathbb{Z}},$
satisfying
(4.7) $\begin{split}k,\widetilde{k}\leqslant K,\quad m,\widetilde{m}\leqslant M,&\quad n,\widetilde{n}\leqslant N,\quad u,\widetilde{u}\sim U,\quad|t|\leqslant T,\\ (ak+bmn)\widetilde{n}\widetilde{u}&=(a\widetilde{k}+b\widetilde{m}\widetilde{n})nu+tp,\\ (ak+bmn)&(a\widetilde{k}+b\widetilde{m}\widetilde{n})\neq 0,\end{split}$
with
(4.8) $T=1+(|a|K+|b|MN)NU/p.$
Note that $\widetilde{n}\mid a\widetilde{k}un+tp$ and $a\widetilde{k}un\not\equiv 0\pmod{p}$. Hence $a\widetilde{k}un+tp\neq 0$ for any $t\in{\mathbb{Z}}$ and thus, for given $\widetilde{k},n,t,u$, which we can fix in $O\left(KNTU\right)$ ways the number of $\widetilde{n}$ satisfying (4.7) is at most $p^{o(1)}$. After $\widetilde{k},\widetilde{m},n,\widetilde{n},t,u$ are fixed, there are at most $(K/N+1)p^{o(1)}$ possibilities for the tuple $(m,\widetilde{u},k)$. This is due to the observations that $\widetilde{u}\mid(a\widetilde{k}+b\widetilde{m}\widetilde{n})un+tp\neq 0$ and the number of solutions $(m,k)$ to
$ak+bmn=\frac{(a\widetilde{k}+b\widetilde{m}\widetilde{n})un+tp}{\widetilde{u}\widetilde{n}}$
with $m\leqslant M,k\leqslant K$ is at most $O\left(1+K/N\right)$ since $k$ falls in a prescribed arithmetic progression modulo $n\sim N$, which is already fixed.
Taking all possibilities into account, we find the contributions from $ak+bmn\neq 0$ to ${\mathcal{I}}_{2}$ are at most
$KNTU\cdot M\cdot\left(1+\frac{K}{N}\right)p^{o(1)}\leqslant KMTU(K+N)p^{o(1)}.$
[…]
To bound ${\mathcal{I}}_{3}$, we apply Lemma 3.1 with ${\mathcal{A}}=\{bv:v\sim V\}$, getting
(4.9) ${\mathcal{I}}_{3}\ll V^{2r}p^{\frac{1}{2}}+V^{r}p$
as a consequence of Weil’s bound for complete character sums over finite fields (see [IK04, Theorem 11.23]).
4.4. Concluding the proof of Theorem 2.1
To balance the two terms in (4.9), we choose
$V=p^{\frac{1}{2r}}\qquad\mbox{and}\qquad U=\frac{1}{4}Mp^{-\frac{1}{2r}}$
in view of (4.2), so that (4.5) and(4.6) become
${\mathcal{I}}_{1}\ll KM^{2}Np^{-\frac{1}{2r}}\qquad\mbox{and}\qquad{\mathcal{I}}_{2}\leqslant KM^{2}p^{-\frac{1}{2r}}(K+N)(1+\mathscr{L}_{1}p^{-1-\frac{1}{2r}})p^{o(1)}$
respectively, where $\mathscr{L}_{1}$ is defined by (2.1), while (4.9) becomes
${\mathcal{I}}_{3}\ll p^{\frac{3}{2}}.$
Now Theorem 2.1 follows by combining the inequality (4.4) and the above estimates for ${\mathcal{I}}_{1}$, ${\mathcal{I}}_{2}$ and ${\mathcal{I}}_{3}$.
5. Proof of Theorem 2.5
5.1. Cauchy–Schwarz inequality and amplification
We note that the auxiliary parameters $T,U,V$ from § 4 have different meaning henceforth. This time it is convenient to assume that $\|\bm{\alpha}\|_{\infty}\leqslant 1$. By the Cauchy–Schwarz inequality, we have
(5.1) $|\mathfrak{T}(\bm{\alpha},\bm{\beta})|^{2}\leqslant\|\bm{\beta}\|_{2}^{2}T,$
where
$T=\sum_{k\leqslant K}\sum_{n\in{\mathbb{Z}}}W\left(\frac{n}{N}\right)\left|\sum_{m\leqslant M}\alpha_{m}\chi(ak+bmn)\right|^{2}$
with any fixed real-valued smooth function $W\in{\mathcal{C}}_{c}^{\infty}([-2,2])$, which majorates the characteristic function of the unit interval $[0,1]$.
Squaring out, changing the order of summation and estimating the contribution from the diagonal terms with $m_{1}=m_{2}$, we obtain
(5.2) $T\ll T_{1}+KMN$
with
$T_{1}=\sum_{k\leqslant K}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\leqslant M\\ m_{1}\neq m_{2}\end{subarray}}\sum_{n\in{\mathbb{Z}}}W\left(\frac{n}{N}\right)\alpha_{m_{1}}\overline{\alpha}_{m_{2}}\chi(ak+bm_{1}n)\overline{\chi}(ak+bm_{2}n).$
For any integers $u,v$ we may write
$T_{1}=\sum_{k\leqslant K}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\leqslant M\\ m_{1}\neq m_{2}\end{subarray}}\sum_{n\in{\mathbb{Z}}}W\left(\frac{n+uv}{N}\right)\alpha_{m_{1}}\overline{\alpha}_{m_{2}}\chi(ak+bm_{1}(n+uv))\overline{\chi}(ak+bm_{2}(n+uv)).$
By Fourier inversion and summing over $u\sim U,v\sim V$ for some positive parameters $U$ and $V$ with
(5.3) $UV=\frac{1}{4}N,$
it follows that
$T_{1}$ $\ll\frac{1}{N}\sum_{k\leqslant K}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\leqslant M\\ m_{1}\neq m_{2}\end{subarray}}\sum_{|n|\leqslant 2N}\sum_{u\sim U}\int_{\mathbb{R}}|\widehat{W}(\xi)|$
×$\qquad\qquad\times\biggl{|}\sum_{v\sim V}\mathrm{e}\left(\frac{uv\xi}{N}\right)\chi(\overline{m_{1}u}(ak+bm_{1}n)+bv)\overline{\chi}(\overline{m_{2}u}(ak+bm_{2}n)+bv)\biggr{|}\mathrm{d}\xi.$
In view of
$\widehat{W}(\xi)\ll(1+|\xi|)^{-2}$
implied by the smoothness of $W(x)$ via partial integration, and making the change of variable $\xi\rightarrow\xi/u$, after simple transformations, we obtain
$T_{1}$ $\ll\frac{1}{NU}\int_{\mathbb{R}}\left(1+\frac{|\xi|}{U}\right)^{-2}\sum_{k\leqslant K}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\leqslant M\\ m_{1}\neq m_{2}\end{subarray}}\sum_{|n|\leqslant 2N}\sum_{u\sim U}$
×$\qquad\qquad\times\biggl{|}\sum_{v\sim V}\mathrm{e}\left(\frac{v\xi}{N}\right)\chi(\overline{m_{1}u}(ak+bm_{1}n)+bv)\overline{\chi}(\overline{m_{2}u}(ak+bm_{2}n)+bv)\biggr{|}\mathrm{d}\xi.$
Again, as in the proof of Theorem 2.1, this implies
(5.4) $\begin{split}T_{1}&\ll\frac{1}{N}\sum_{k\leqslant K}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\leqslant M\\ m_{1}\neq m_{2}\end{subarray}}\sum_{|n|\leqslant 2N}\sum_{u\sim U}\\ &\qquad\quad\times\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(\overline{m_{1}u}(ak+bm_{1}n)+bv)\overline{\chi}(\overline{m_{2}u}(ak+bm_{2}n)+bv)\right|\end{split}$
for some $\xi\in{\mathbb{R}}$.
5.2. Regrouping, counting and Weil’s bound
Put
$\varrho(x_{1},x_{2})=\mathop{\sum_{|n|\leqslant 2N}\sum_{k\leqslant K}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\sum_{u\sim U}}_{\begin{subarray}{c}ak+bm_{1}n\equiv um_{1}x_{1}\not\equiv 0\bmod p\\ ak+bm_{1}n\equiv um_{2}x_{2}\not\equiv 0\bmod p\\ m_{1}\neq m_{2}\end{subarray}}1$
for $x_{1},x_{2}\bmod p$. We now have
$T_{1}\ll\frac{1}{N}\mathop{\sum\sum}_{x_{1},x_{2}\bmod p}\varrho(x_{1},x_{2})\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(x_{1}+bv)\overline{\chi}(x_{2}+bv)\right|+(1+MN/p)KMp^{o(1)},$
where the last term comes from those $k,m_{1},m_{2},n$ with
$p\mid(ak+bm_{1}n)(ak+bm_{2}n)$
in (5.4).
Applying the Hölder inequality with $r\geqslant 1$ similarly to (4.4), we find
(5.5) $T_{1}\ll\frac{1}{N}{\mathcal{J}}_{1}^{1-\frac{1}{r}}({\mathcal{J}}_{2}\cdot{\mathcal{J}}_{3})^{\frac{1}{2r}}+(1+MN/p)KMp^{o(1)}$
with
${\mathcal{J}}_{j}=\mathop{\sum\sum}_{x_{1},x_{2}\bmod p}\varrho(x_{1},x_{2})^{j},\quad j=1,2,$
and
${\mathcal{J}}_{3}=\mathop{\sum\sum}_{x_{1},x_{2}\bmod p}\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(x_{1}+bv)\overline{\chi}(x_{2}+bv)\right|^{2r}.$
Note that
(5.6) ${\mathcal{J}}_{1}\ll KM^{2}NU.$
We appeal to the following lemma in bounding ${\mathcal{J}}_{2}$.
Lemma 5.1.
With the above notation, we have
(5.7) ${\mathcal{J}}_{2}\leqslant KNU(K+M^{2})\left(1+\frac{\left(|a|K+|b|MN\right)MU}{p}\right)^{2}p^{o(1)}.$
Proof.
We observe that ${\mathcal{J}}_{2}$ is bounded by the number of 10-tuples
$(k,\widetilde{k},m_{1},\widetilde{m}_{1},m_{2},\widetilde{n}_{2},n,\widetilde{n},u,\widetilde{u})\in{\mathbb{N}}^{10}$
satisfying
$k,\widetilde{k}\leqslant K,\quad m_{1},\widetilde{m}_{1},m_{2},\widetilde{m}_{2}$ $\leqslant M,\quad n,\widetilde{n}\in[-2N,2N],\quad u,\widetilde{u}\sim U,$
$(ak+bm_{1}n)\widetilde{m}_{1}\widetilde{u}$ $\equiv(a\widetilde{k}+b\widetilde{m}_{1}\widetilde{n})m_{1}u\not\equiv 0\bmod p,$
$(ak+bm_{2}n)\widetilde{m}_{2}\widetilde{u}$ $\equiv(a\widetilde{k}+b\widetilde{m}_{2}\widetilde{n})m_{2}u\not\equiv 0\bmod p,$
$m_{1}$ $\neq m_{2},\quad\widetilde{m}_{1}\neq\widetilde{m}_{2}.$
We now rewrite the above congruences as equations
(5.8a) $(ak+bm_{1}n)\widetilde{m}_{1}\widetilde{u}$ $=(a\widetilde{k}+b\widetilde{m}_{1}\widetilde{n})m_{1}u+t_{1}p\neq 0,$
(5.8b) $(ak+bm_{2}n)\widetilde{m}_{2}\widetilde{u}$ $=(a\widetilde{k}+b\widetilde{m}_{2}\widetilde{n})m_{2}u+t_{2}p\neq 0,$
where $0\leqslant|t_{1}|,|t_{2}|\leqslant T$ with
(5.9) $T=1+(|a|K+2|b|MN)MU/p.$
It follows from (5.8a) that $\widetilde{m}_{1}\mid a\widetilde{k}m_{1}u+t_{1}p\neq 0$ and $\widetilde{u}\mid(\widetilde{m}_{1}\widetilde{n}+a\widetilde{k})m_{1}u+t_{1}p\neq 0$, which produces at most $p^{o(1)}$ tuples of $(\widetilde{m}_{1},\widetilde{u})$ for given $\widetilde{n},\widetilde{k},m_{1},u,t_{1}$. After fixing $\widetilde{m}_{1},\widetilde{u},\widetilde{n},\widetilde{k},m_{1},u,t_{1}$, we obtain the equation
(5.10) $ak+bm_{1}n=\frac{(a\widetilde{k}+b\widetilde{m}_{1}\widetilde{n})m_{1}u+t_{1}p}{\widetilde{m}_{1}\widetilde{u}}$
in $k$ and $n$, thanks to (5.8a). We further fix $m_{2}$ and $t_{2}$. We see from (5.8b) that $\widetilde{m}_{2}\mid a\widetilde{k}m_{2}u+t_{2}p\neq 0$, by which there are at most $p^{o(1)}$ values of $\widetilde{m}_{2}$ for given $\widetilde{k},m_{2},u,t_{2}$. As before, using (5.8b), we obtain the equation
(5.11) $ak+bm_{2}n=\frac{(a\widetilde{k}+b\widetilde{m}_{2}\widetilde{n})m_{2}u+t_{2}p}{\widetilde{m}_{2}\widetilde{u}}$
in $k$ and $n$. It then follows from (5.10) and (5.11) that $k$ is uniquely defined modulo $m_{1}$ and modulo $m_{2}$, so that the number of such positive integers $k$ is at most
$\frac{K}{\operatorname{lcm}[m_{1},m_{2}]}+1=\frac{\gcd(m_{1},m_{2})}{m_{1}m_{2}}K+1.$
Now $n$ is uniquely determined after $k$ is fixed. Collecting all above arguments, we find
${\mathcal{J}}_{2}$ $\leqslant KNUT^{2}p^{o(1)}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\left(\frac{\gcd(m_{1},m_{2})}{m_{1}m_{2}}K+1\right)$
$\leqslant KNUT^{2}(K+M^{2})p^{o(1)}.$
[…]
Using Lemma 3.2 with the special choice ${\mathcal{A}}=\{bv:v\sim V\}$ therein, we find
(5.12) ${\mathcal{J}}_{3}\ll V^{2r}p+V^{r}p^{2}.$
5.3. Concluding the proof of Theorem 2.5
We now choose
$V=p^{\frac{1}{r}}\qquad\mbox{and}\qquad U=\frac{1}{4}Np^{-\frac{1}{r}}$
subject to the constraint in (5.3), so that the above bounds for ${\mathcal{J}}_{1}$ and ${\mathcal{J}}_{2}$ in (5.6) and (5.7) become
${\mathcal{J}}_{1}\ll K(MN)^{2}p^{-\frac{1}{r}}\qquad\mbox{and}\qquad{\mathcal{J}}_{2}\leqslant KN^{2}p^{-\frac{1}{r}}(K+M^{2})(1+\mathscr{L}p^{-1-\frac{1}{r}})^{2}p^{o(1)},$
respectively, while the bound for ${\mathcal{J}}_{3}$ in  (5.12) becomes
${\mathcal{J}}_{3}\ll p^{3}.$
Substituting these estimates to (5.5), we arrive that
$T_{1}\leqslant\frac{1}{N}(KM^{2}N^{2})^{1-\frac{1}{2r}}(1+KM^{-2})^{\frac{1}{2r}}$ $\left(1+\mathscr{L}_{1}p^{-1-\frac{1}{r}}\right)^{\frac{1}{r}}$
×$\qquad\times p^{\frac{1}{2r}+\frac{1}{2r^{2}}+o(1)}+(1+MN/p)KMp^{o(1)},$
from which and then inequalities (5.1) and (5.2), Theorem 2.5 follows.
6. Proof of Theorem 2.7
6.1. Preliminary transformations
Assume $\|\bm{\alpha}\|_{\infty}\leqslant 1$. By the Cauchy–Schwarz inequality, we have
(6.1) $|\mathfrak{Q}(\bm{\alpha},\bm{\beta})|^{2}\leqslant\|\bm{\beta}\|_{2}^{2}Q,$
where
$Q=\mathop{\sum\sum}_{k,n\in{\mathbb{Z}}}\varPhi\left(\frac{k}{K}\right)\varPhi\left(\frac{n}{N}\right)\left|\sum_{\ell\leqslant L}\sum_{m\leqslant M}\alpha_{\ell,m}\chi(ak\ell+bmn)\right|^{2},$
for any smooth function $\varPhi$, which dominates the characteristic function of $[-1,1]$ and supported only inside of the interval $[-2,2]$. Squaring out and switching the order of summation, we get
(6.2) $Q\leqslant Q_{1}+KLMN(LM/p+1)p^{o(1)}$
with
$Q_{1}=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}\mathop{\sum\sum}_{k,n\in{\mathbb{Z}}}\varPhi\left(\frac{k}{K}\right)$ $\varPhi\left(\frac{n}{N}\right)\alpha_{\ell_{1},m_{1}}\overline{\alpha}_{\ell_{2},m_{2}}$
×$\quad\times\chi(ak\ell_{1}+bm_{1}n)\overline{\chi}(ak\ell_{2}+bm_{2}n),$
where the second term in (6.2) comes from the contribution from $\ell_{1}m_{2}\equiv\ell_{2}m_{1}\bmod p$. The trivial bound for $Q_{1}$ is
(6.3) $Q_{1}\ll KL^{2}M^{2}N,$
and, by (6.1), we obtain a non-trivial bound of $\mathfrak{Q}(\bm{\alpha},\bm{\beta})$ as soon as we improve (6.3). For all integers $(u,v,w)$, using that ${\mathbb{Z}}$ is invariant under shifts by integers, we can write
$Q_{1}$ $=\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}\mathop{\sum\sum}_{k,n\in{\mathbb{Z}}}\varPhi\left(\frac{k+wv}{K}\right)\varPhi\left(\frac{n+uv}{N}\right)\alpha_{\ell_{1},m_{1}}\overline{\alpha}_{\ell_{2},m_{2}}$
×$\qquad\qquad\quad\times\chi\left(a(k+uv)\ell_{1}+bm_{1}(n+wv)\right)\overline{\chi}\left(a(k+uv)\ell_{2}+bm_{2}(n+wv)\right).$
By Fourier inversion and summing over $u\sim U,v\sim V,w\sim W$ with
(6.4) $UV=\frac{1}{4}K\qquad\mbox{and}\qquad WV=\frac{1}{4}N,$
it follows that
(6.5) $\begin{split}Q_{1}&\ll\frac{1}{UVW}\sum_{|k|\leqslant 2K}\sum_{|n|\leqslant 2N}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}\sum_{u\sim U}\sum_{w\sim W}\iint_{{\mathbb{R}}^{2}}|\widehat{\varPhi}(\xi)\widehat{\varPhi}(\eta)|\\ &\qquad\qquad\quad\times\biggl{|}\sum_{v\sim V}\mathrm{e}\left(\frac{uv\xi}{N}+\frac{wv\eta}{K}\right)\chi\left(\overline{(a\ell_{1}u+bm_{1}w)}(ak\ell_{1}+bm_{1}n)+v\right)\\ &\qquad\qquad\qquad\qquad\quad\times\overline{\chi}\left(\overline{(a\ell_{2}u+bm_{2}w)}(ak\ell_{2}+bm_{2}n)+v\right)\biggr{|}\mathrm{d}\xi\mathrm{d}\eta+\mathsf{Err}_{1},\end{split}$
where
  • the variables of summation satisfy the extra condition
    (6.6) $p\nmid(a\ell_{1}u+bm_{1}w)(a\ell_{2}u+bm_{2}w)(ak\ell_{1}+bm_{1}n)(ak\ell_{2}+bm_{2}n),$
  • the term $\mathsf{Err}_{1}$ corresponds to the error induced by the terms
    $p\mid(a\ell_{1}u+bm_{1}w)(a\ell_{2}u+bm_{2}w)(ak\ell_{1}+bm_{1}n)(ak\ell_{2}+bm_{2}n).$
As a typical possibility, the number of tuples $(\ell_{1},u,m_{1},w)$ satisfying $p\mid(a\ell_{1}u+bm_{1}w)$ is at most $LU(1+MW/p)p^{o(1)}$. Taking all the remaining possibilities into account, we find
(6.7) $\mathsf{Err}_{1}\ll\frac{KL^{2}MNU}{UVW}\left(1+\frac{MW}{p}\right)p^{o(1)}\ll KL^{2}M\left(1+\frac{MN}{pV}\right)p^{o(1)}.$
Combining (6.7) with the inequality $\widehat{\varPhi}(\xi)\ll(1+|\xi|)^{-2}$ and with the change of variables $(\xi,\eta)\rightarrow(\xi/u,\eta/w)$ in (6.5), we obtain
$Q_{1}$ $\ll\frac{1}{U^{2}VW^{2}}\iint_{{\mathbb{R}}^{2}}\left(1+\frac{|\xi|}{U}\right)^{-2}\left(1+\frac{|\eta|}{W}\right)^{-2}\sum_{|k|\leqslant 2K}\sum_{|n|\leqslant 2N}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}$
×$\qquad\qquad\qquad\times\sum_{u\sim U}\sum_{w\sim W}\biggl{|}\sum_{v\sim V}\mathrm{e}\left(\frac{v\xi}{N}+\frac{v\eta}{K}\right)\chi\left(\overline{(a\ell_{1}u+bm_{1}w)}(ak\ell_{1}+bm_{1}n)+v\right)$
×$\qquad\qquad\qquad\qquad\qquad\quad\qquad\times\overline{\chi}\left(\overline{(a\ell_{2}u+bm_{2}w)}(ak\ell_{2}+bm_{2}n)+v\right)\biggr{|}\mathrm{d}\xi\mathrm{d}\eta\,+\mathsf{Err}_{1}.$
This implies
$Q_{1}\ll\frac{1}{UVW}\sum_{|k|\leqslant 2K}$ $\sum_{|n|\leqslant 2N}\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}$
×$\quad\times\sum_{u\sim U}\sum_{w\sim W}\Bigl{|}\sum_{v\sim V}\mathrm{e}(v\xi)\chi\left(\overline{(a\ell_{1}u+bm_{1}w)}(ak\ell_{1}+bm_{1}n)+v\right)$
×$\qquad\qquad\quad\quad\times\overline{\chi}\left(\overline{(a\ell_{2}u+bm_{2}w)}(ak\ell_{2}+bm_{2}n)+v\right)\Bigr{|}+\mathsf{Err}_{1}$
for some $\xi\in{\mathbb{R}}$, where $\mathsf{Err}_{1}$ satisfies (6.7). We now pull out $d=\gcd(ak,bn)$, for which we have $d\leqslant D:=\min\{|a|K,|b|N\}$. Hence, changing the variables $k\mapsto dk$ and $n\mapsto dn$, we obtain the inequality
(6.8) $Q_{1}\ll\frac{1}{UVW}\sum_{d\leqslant D}Q_{1}(d)\,+\mathsf{Err}_{1}$
with
$Q_{1}(d)=\mathop{\sum_{|k|\leqslant 2K/d}\sum_{|n|\leqslant 2N/d}}_{\gcd(ak,bn)=1}$ $\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}\ell_{1},\ell_{2}\leqslant L,\leavevmode\nobreak\ m_{1},m_{2}\leqslant M\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p\end{subarray}}$
×$\times\sum_{u\sim U}\sum_{w\sim W}\biggl{|}\sum_{v\sim V}\mathrm{e}(v\xi)\chi\left(d\overline{(a\ell_{1}u+bm_{1}w)}(ak\ell_{1}+bm_{1}n)+v\right)$
×$\qquad\qquad\qquad\qquad\times\overline{\chi}\left(d\overline{(a\ell_{2}u+m_{2}w)}(ak\ell_{2}+bm_{2}n)+v\right)\biggr{|}.$
Of course the variables of summation continue to satisfy (6.6). We further impose a restriction that $kw\neq nu$, which introduces an additional error $\mathsf{Err}_{2}(d)$ with
$\mathsf{Err}_{2}(d)\ll KL^{2}M^{2}VWd^{-1}p^{o(1)}.$
Next for $x_{1},x_{2}\in{\mathbb{F}}_{p}^{*}$, $1\leqslant A\ll\mathscr{L}_{2}/d$, $1\leqslant B\ll|a|LU+|b|MW\ll\mathscr{L}_{2}/(dV)$ and $1\leqslant C\leqslant K/d$, we put
$\varrho(x_{1},x_{2})=\varrho(x_{1},x_{2};A,B,C)=\mathop{\sum_{|k|\sim C}\sum_{|n|\leqslant 2N/d}\mathop{\sum\sum}_{\ell_{1},\ell_{2}\leqslant L}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\sum_{u\sim U}\sum_{w\sim W}}_{\begin{subarray}{c}ak\ell_{1}+bm_{1}n\equiv(a\ell_{1}u+bm_{1}w)x_{1}\bmod p\\ ak\ell_{2}+bm_{2}n\equiv(a\ell_{2}u+bm_{2}w)x_{2}\bmod p\\ \ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p,\quad kw\neq nu,\quad\gcd(ak,bn)=1\\ |ak\ell_{1}+bm_{1}n|\sim A,\leavevmode\nobreak\ |a\ell_{2}u+bm_{2}w|\sim B\end{subarray}}1,$
where the condition (6.6) continues to apply to the variables of summation. We now have
$Q_{1}(d)\leqslant p^{o(1)}\sup_{\begin{subarray}{c}1\leqslant A\ll\mathscr{L}_{2}/d\\ 1\leqslant B\ll\mathscr{L}_{2}/(dV)\\ 1\leqslant C\leqslant K/d\end{subarray}}\sum_{x_{1},x_{2}\in{\mathbb{F}}_{p}^{*}}\varrho(x_{1},x_{2})\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(x_{1}+v)\overline{\chi}(x_{2}+v)\right|+\mathsf{Err}_{2}(d).$
Applying the Hölder inequality with $r\in{\mathbb{N}}$, we find
(6.9) $Q_{1}(d)\leqslant Q_{2}(d)\,p^{o(1)}+\mathsf{Err}_{2}(d)$
where
(6.10) $Q_{2}(d)=\sup_{\begin{subarray}{c}1\leqslant A\ll\mathscr{L}_{2}/d\\ 1\leqslant B\ll\mathscr{L}_{2}/(dV)\\ 1\leqslant C\leqslant K/d\end{subarray}}\Sigma_{1}^{1-\frac{1}{r}}\left(\Sigma_{2}\cdot\Sigma_{3}\right)^{\frac{1}{2r}}$
with
$\Sigma_{j}=\Sigma_{j}(A,B,C)=\mathop{\sum\sum}_{x_{1},x_{2}\in{\mathbb{F}}_{p}^{*}}\varrho(x_{1},x_{2};A,B,C)^{j},\qquad j=1,2,$
and
$\Sigma_{3}=\mathop{\sum\sum}_{x_{1},x_{2}\in{\mathbb{F}}_{p}}\left|\sum_{v\sim V}\mathrm{e}(v\xi)\chi(x_{1}+v)\overline{\chi}(x_{2}+v)\right|^{2r}.$
Trivially we have
(6.11) $\Sigma_{1}\ll CL^{2}M^{2}NUWd^{-1}.$
Furthermore, using Lemma 3.2, we have the inequality
(6.12) $\Sigma_{3}\ll V^{2r}p+V^{r}p^{2}.$
Next we estimate $\Sigma_{2}$.
6.2. Bounding $\Sigma_{2}$
This following bound is the core of our method.
Lemma 6.1.
With the above notation, we have
$\Sigma_{2}\leqslant(M/C+1)^{2}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{2}(ABLM)^{1+o(1)}(CW+NU/d)$
for all $1\leqslant A\ll\mathscr{L}_{2}/d$, $1\leqslant B\ll\mathscr{L}_{2}/(dV)$ and $1\leqslant C\leqslant K/d$.
Proof.
Note that $\Sigma_{2}$ is bounded by the number of tuples
$(k,\widetilde{k},\ell_{1},\widetilde{\ell}_{1},\ell_{2},\widetilde{\ell}_{2},m_{1},\widetilde{m}_{1},m_{2},\widetilde{m}_{2},n,\widetilde{n},u,\widetilde{u},w,\widetilde{w})$
satisfying
$|k|,|\widetilde{k}|\sim C,$ $\quad\ell_{1},\widetilde{\ell}_{1},\ell_{2},\widetilde{\ell}_{2}\leqslant L,\quad m_{1},\widetilde{m}_{1},m_{2},\widetilde{m}_{2}\leqslant M,\quad|n|,|\widetilde{n}|\leqslant 2N/d,$
$u,\widetilde{u}\sim U,$ $\quad w,\widetilde{w}\sim W,\quad kw\neq nu,\quad\widetilde{k}\widetilde{w}\neq\widetilde{n}\widetilde{u},\quad\gcd(ak,bn)=\gcd(a\widetilde{k},b\widetilde{n})=1,$
$|ak\ell_{1}+bm_{1}n|$ $\sim A,\quad|a\widetilde{k}\widetilde{\ell}_{1}+b\widetilde{m}_{1}\widetilde{n}|\sim A,\quad|a\ell_{2}u+bm_{2}w|\sim B,\quad|a\widetilde{\ell}_{2}\widetilde{u}+b\widetilde{m}_{2}\widetilde{w}|\sim B,$
$\qquad\ell_{1}m_{2}\not\equiv\ell_{2}m_{1}\bmod p,\qquad\widetilde{\ell}_{1}\widetilde{m}_{2}\not\equiv\widetilde{\ell}_{2}\widetilde{m}_{1}\bmod p,$
$p\nmid$ $(a\ell_{1}u+bm_{1}w)(a\ell_{2}u+bm_{2}w)(ak\ell_{1}+bm_{1}n)(ak\ell_{2}+bm_{2}n),$
$p\nmid$ $(a\widetilde{\ell}_{1}\widetilde{u}+b\widetilde{m}_{1}\widetilde{w})(a\widetilde{\ell}_{2}\widetilde{u}+b\widetilde{m}_{2}\widetilde{w})(a\widetilde{k}\widetilde{\ell}_{1}+b\widetilde{m}_{1}\widetilde{n})(a\widetilde{k}\widetilde{\ell}_{2}+b\widetilde{m}_{2}\widetilde{n}),$
and also such that
(6.13) $(ak\ell_{1}+bm_{1}n)(a\widetilde{\ell}_{1}\widetilde{u}+b\widetilde{m}_{1}\widetilde{w})=(a\widetilde{k}\widetilde{\ell}_{1}+b\widetilde{m}_{1}\widetilde{n})(a\ell_{1}u+bm_{1}w)+z_{1}p,$
and
(6.14) $(ak\ell_{2}+bm_{2}n)(a\widetilde{\ell}_{2}\widetilde{u}+b\widetilde{m}_{2}\widetilde{w})=(a\widetilde{k}\widetilde{\ell}_{2}+b\widetilde{m}_{2}\widetilde{n})(a\ell_{2}u+bm_{2}w)+z_{2}p,$
with some $0\leqslant|z_{1}|,|z_{2}|\leqslant Z$ where $Z=1+2(|a|KL+|b|MN)(|a|LU+|b|MW)/(dp)$, so by hypothesis we have the inequality
(6.15) $Z\ll 1+\frac{\mathscr{L}_{2}^{2}}{dpV}.$
We now fix $m_{1},m_{2},n,k,\ell_{1},\ell_{2},u,w$, and put
$a_{j}=ak\ell_{j}+bm_{j}n,\quad b_{j}=a\ell_{j}u+bm_{j}w$
for $j=1,2$. Note that
$|a_{1}|\sim A,\leavevmode\nobreak\ |a_{2}|\leqslant 2\mathscr{L}_{2}/d,\leavevmode\nobreak\ |b_{1}|\leqslant\mathscr{L}_{2}/(dV)\text{ and }|b_{2}|\sim B$
with $a_{1}a_{2}b_{1}b_{2}\neq 0$. Given $z_{1},z_{2}$ with $0\leqslant|z_{1}|,|z_{2}|\leqslant Z$ as above, we now look at the equations
(6.16) $a_{1}x_{1}=b_{1}y_{1}+z_{1}p,\quad a_{2}x_{2}=b_{2}y_{2}+z_{2}p$
in $x_{1},x_{2},y_{1},y_{2}\in{\mathbb{Z}}$ with $|x_{1}|\leqslant\mathscr{L}_{2}/(dV)$, $|x_{2}|\sim B$, $|y_{1}|\sim A$ and $|y_{2}|\leqslant 2\mathscr{L}_{2}/d$. It is clear that the number $Q$ of such tuples $(x_{1},x_{2},y_{1},y_{2})$ satisfying (6.16) is
(6.17) $Q\ll\left(\frac{A\gcd(a_{1},b_{1})}{a_{1}}+1\right)\left(\frac{B\gcd(a_{2},b_{2})}{b_{2}}+1\right)\ll\gcd(a_{1},b_{1})\gcd(a_{2},b_{2}).$
We now fix one of such tuples $(x_{1},x_{2},y_{1},y_{2})$, satisfying (6.16). Then (6.13) and (6.14) lead to the system of four equations
(6.18) $\begin{cases}a\widetilde{\ell}_{1}\widetilde{u}+b\widetilde{m}_{1}\widetilde{w}=x_{1},\\ a\widetilde{k}\widetilde{\ell}_{1}+b\widetilde{m}_{1}\widetilde{n}=y_{1},\\ a\widetilde{\ell}_{2}\widetilde{u}+b\widetilde{m}_{2}\widetilde{w}=x_{2},\\ a\widetilde{k}\widetilde{\ell}_{2}+b\widetilde{m}_{2}\widetilde{n}=y_{2},\end{cases}$
in $\widetilde{k},\widetilde{\ell}_{1},\widetilde{\ell}_{2},\widetilde{m}_{1},\widetilde{m}_{2},\widetilde{n},\widetilde{u},\widetilde{w}\in{\mathbb{Z}}$. Recalling the conditions ~$\widetilde{\ell}_{1}\widetilde{m}_{2}\not\equiv\widetilde{\ell}_{2}\widetilde{m}_{1}\bmod p$, it suffices to consider the solutions satisfying ~$\widetilde{\ell}_{1}\widetilde{m}_{2}\neq\widetilde{\ell}_{2}\widetilde{m}_{1}$.
We now fix integers $\mathfrak{l}_{1},\mathfrak{l}_{2},\mathfrak{m}_{1},\mathfrak{m}_{2}$ satisfying
(6.19) $\mathfrak{l}_{1},\mathfrak{l}_{2}\leqslant L,\quad\mathfrak{m}_{1},\mathfrak{m}_{2}\leqslant M,\quad\mathfrak{l}_{1}\mathfrak{m}_{2}\neq\mathfrak{l}_{2}\mathfrak{m}_{1}.$
such that the two equations
(6.20) $a\widetilde{k}\mathfrak{l}_{1}+b\mathfrak{m}_{1}\widetilde{n}=y_{1},\quad a\widetilde{k}\mathfrak{l}_{2}+b\mathfrak{m}_{2}\widetilde{n}=y_{2}$
are solvable in $\widetilde{k},\widetilde{n}$ with $\gcd(\widetilde{k},\widetilde{n})=1$, given the above $\mathfrak{l}_{1},\mathfrak{l}_{2},\mathfrak{m}_{1},\mathfrak{m}_{2}$. Note that the system (6.20) has at most one solution $(\widetilde{k},\widetilde{n})$ because its determinant does not vanish. Suppose that a solution to (6.20) does exist. Then by (6.19) all solutions to the second equation in (6.18) are of the shape
~$\widetilde{\ell}_{1}=\mathfrak{l}_{1}-s_{1}b\widetilde{n},\quad\widetilde{m}_{1}=\mathfrak{m}_{1}+s_{1}a\widetilde{k},\quad 0\leqslant|s_{1}|\ll M/|\widetilde{k}|\ll M/C,$
and all solutions to the fourth equation in (6.18) are of the shape
~$\widetilde{\ell}_{2}=\mathfrak{l}_{2}-s_{2}b\widetilde{n},\quad\widetilde{m}_{2}=\mathfrak{m}_{2}+s_{2}a\widetilde{k},\quad 0\leqslant|s_{2}|\ll M/|\widetilde{k}|\ll M/C.$
After determining ~$\widetilde{\ell}_{1},\widetilde{\ell}_{2},\widetilde{m}_{1},\widetilde{m}_{2}$ with ~$\widetilde{\ell}_{1}\widetilde{m}_{2}\neq\widetilde{\ell}_{2}\widetilde{m}_{1}$, we may find at most one tuple $(\widetilde{u},\widetilde{w})$ satisfying the first and third equations in (6.18) simultaneously.
Combining all above arguments, in particular, recalling (6.17), we derive that
$\Sigma_{2}$ $\ll(M/C+1)^{2}Z^{2}\mathop{\sum_{|k|\sim C}\mathop{\sum\sum}_{\ell_{1},\ell_{2}\leqslant L}\mathop{\sum\sum}_{m_{1},m_{2}\leqslant M}\sum_{|n|\leqslant 2N/d}\sum_{u\sim U}\sum_{w\sim W}}_{\begin{subarray}{c}ak\ell_{1}+bm_{1}n=a_{1},\leavevmode\nobreak\ a\ell_{1}u+bm_{1}w=b_{1}\\ ak\ell_{2}+bm_{2}n=a_{2},\leavevmode\nobreak\ a\ell_{2}u+bm_{2}w=b_{2}\\ |a_{1}|\sim A,\leavevmode\nobreak\ |b_{2}|\sim B,\leavevmode\nobreak\ kw\neq nu\end{subarray}}\gcd(a_{1},b_{1})\gcd(a_{2},b_{2}).$
[…]
Subsequently, by Lemma 6.1, for all $1\leqslant A\ll\mathscr{L}_{2}/d$, $1\leqslant B\ll|a|LU+|b|MW\ll\mathscr{L}_{2}/(dV)$ and $1\leqslant C\leqslant K/d$, we have the bound
(6.21) $\Sigma_{2}\leqslant(M/C+1)^{2}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{2}(\mathscr{L}_{2}^{2}LM)^{1+o(1)}(CW+NU/d)d^{-2}V^{-1}.$
6.3. Concluding the proof of Theorem 2.7
We substitute the bounds (6.11), (6.12) and (6.21) into (6.10) and derive
$Q_{2}(d)$ $\leqslant p^{o(1)}\sup_{\begin{subarray}{c}C\leqslant K/d\end{subarray}}\left(CL^{2}M^{2}NUWd^{-1}\right)^{1-\frac{1}{r}}(M/C+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{\frac{1}{r}}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\mathscr{L}_{2}^{2}LM(CW+NU/d)d^{-2}V^{-1}\right)^{\frac{1}{2r}}(V^{2r}p+V^{r}p^{2})^{\frac{1}{2r}}$
$\leqslant p^{o(1)}\left(KL^{2}M^{2}NUWd^{-2}\right)^{1-\frac{1}{r}}(dM/K+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{\frac{1}{r}}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\mathscr{L}_{2}^{2}KLMNd^{-3}V^{-2}\right)^{\frac{1}{2r}}(V^{2r}p+V^{r}p^{2})^{\frac{1}{2r}}.$
Recalling (6.8) and (6.9) and then summing over $d\leqslant D$, we obtain
$Q_{1}$ $\ll\frac{p^{o(1)}}{UVW}\sum_{d\leqslant D}\left(Q_{2}(d)+\mathsf{Err}_{2}(d)\right)+\mathsf{Err}_{1}$
$\leqslant p^{o(1)}\left(KL^{2}M^{2}N\right)^{1-\frac{1}{r}}(UVW)^{-\frac{1}{r}}(M/K+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{pV}\right)^{\frac{1}{r}}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\mathscr{L}_{2}^{2}KLMNp\right)^{\frac{1}{2r}}(1+V^{-\frac{1}{2}}p^{\frac{1}{2r}})$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+KL^{2}M^{2}U^{-1}p^{o(1)}+KL^{2}M\left(1+\frac{MN}{pV}\right)p^{o(1)}.$
Taking
$V=p^{\frac{1}{r}},\qquad U=\frac{1}{4}Kp^{-\frac{1}{r}},\qquad W=\frac{1}{4}Np^{-\frac{1}{r}},$
so that the assumption (6.4) is satisfied, we derive that
$Q_{1}$ $\leqslant\left(KL^{2}M^{2}N\right)^{1-\frac{1}{r}}(KN)^{-\frac{1}{r}}(M/K+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{p^{1+\frac{1}{r}}}\right)^{\frac{1}{r}}\left(\mathscr{L}_{2}^{2}KLMN\right)^{\frac{1}{2r}}p^{\frac{1}{2r}+\frac{1}{r^{2}}+o(1)}$
$\qquad\qquad+(LM)^{2}p^{\frac{1}{r}+o(1)}+KL^{2}M\left(1+\frac{MN}{p^{1+\frac{1}{r}}}\right)p^{o(1)}$
$\leqslant\left(KLMN\right)^{2-\frac{3}{2r}}(KN)^{-1}(M/K+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{p^{1+\frac{1}{r}}}\right)^{\frac{1}{r}}\mathscr{L}_{2}^{\frac{1}{r}}p^{\frac{1}{2r}+\frac{1}{r^{2}}+o(1)}$
$\qquad\qquad+(LM)^{2}p^{\frac{1}{r}+o(1)}+KL^{2}M\left(1+\frac{MN}{p^{1+\frac{1}{r}}}\right)p^{o(1)}.$
It remains to recall (6.1) and (6.2), to conclude the proof of Theorem 2.7.
7. Proof of Theorem 2.11
Let $p>\max\{K,M,N\}$. Consider the trilinear sum of Legendre symbols
$\mathfrak{S}=\sum_{k\leqslant K}\sum_{m\in{\mathcal{M}}}\sum_{n\in{\mathcal{N}}}\left(\frac{k+mn}{p}\right)$
with
(7.1) $K=\left\lfloor p^{1+\frac{1}{r}+\eta}(MN)^{-2}+p^{\eta}\right\rfloor.$
Suppose that the values of $k+mn$, for all given $k,m,n$ in the above ranges, are always quadratic non-residues or zero modulo $p$. Then we see from (2.6) that $\mathfrak{S}$ satisfies the trivial bound
$|\mathfrak{S}|\geqslant c_{0}^{2}KMN+O((MN/p+1)Kp^{o(1)}),$
where the error term accounts for the cases with $k+mn\equiv 0\pmod{p}$.
Therefore, Theorem 2.11 is proved as soon as we have the inequality
(7.2) $\mathfrak{S}=o(KMN)$
under the hypotheses (2.5) and $p$ tending to infinity.
We first consider the case $(MN)^{2}>p^{1+\frac{1}{r}}$, for which we see from (7.1) that $K\asymp p^{\eta}$. Now (7.2) is an immediate consequence of Karatsuba presented already in (1.6) and (1.5). We henceforth assume that $(MN)^{2}\leqslant p^{1+\frac{1}{r}}$. Applying Theorem 2.5 with the choice
$a=b=1,\,\alpha_{m}={\bf 1}_{{\mathcal{M}}}(m)\text{ and }\beta_{k,n}={\bf 1}_{[1,K]}(k)\cdot{\bf 1}_{{\mathcal{N}}}(n),$
now (7.2) is proved as soon as one has the following four inequalities
$M\geqslant K^{\frac{1}{2}},\quad(MN)^{2}\leqslant Kp^{1+\frac{1}{r}},\quad K(MN)^{2}\geqslant p^{1+\frac{1}{r}+\frac{\eta}{2}},\quad M\geqslant p^{\frac{\eta}{4}}.$
The choice (7.1) guarantees the above four conditions provided that
$M\geqslant p^{\frac{\eta}{4}}\qquad\mbox{and}\qquad M^{4}N^{2}>p^{1+\frac{1}{r}+\eta}.$
The restriction $M\geqslant p^{\frac{\eta}{4}}$ can also be dropped, since otherwise we should have $N^{2}>p^{1+\frac{1}{r}}$, and we may instead appeal to Karatsuba by noting that $K\geqslant p^{\eta}$ thanks to the choice (7.1). This completes the proof of Theorem 2.11.
Appendix A Karatsuba’s bound for double character sums
Recall that
$\mathfrak{B}(\bm{\alpha},\bm{\beta})=\sum_{m\in{\mathcal{M}}}\sum_{n\in{\mathcal{N}}}\alpha_{m}\beta_{n}\chi(m+n)$
as in (1.3). We now give a proof of (1.6) in the range of (1.5). More precisely, we have the following explicit estimate for $\mathfrak{B}(\bm{\alpha},\bm{\beta})$.
Theorem A.1.
Let $\chi$ be a non-trivial character of ${\mathbb{F}}_{p}^{\times}$ and ${\mathcal{M}},{\mathcal{N}}\subseteq{\mathbb{F}}_{p}$ two arbitrary subsets with cardinalities $M,N$, respectively. For each positive integer $r$, we have
$\mathfrak{B}(\bm{\alpha},\bm{\beta})\ll\|\bm{\alpha}\|_{1}^{1-\frac{1}{r}}\|\bm{\alpha}\|_{2}^{\frac{1}{r}}\|\bm{\beta}\|_{\infty}(Np^{\frac{1}{4r}}+N^{\frac{1}{2}}p^{\frac{1}{2r}}),$
where the implicit constant depends only on $r$.
Remark A.2.
Theorem A.1 implies that
$\mathfrak{B}(\bm{\alpha},\bm{\beta})\ll\|\bm{\alpha}\|_{\infty}\|\bm{\beta}\|_{\infty}MN(M^{-\frac{1}{2r}}p^{\frac{1}{4r}}+M^{-\frac{1}{2r}}N^{-\frac{1}{2}}p^{\frac{1}{2r}}).$
This readily implies (1.6) for
$M>p^{\frac{1}{2}+\eta}\qquad\mbox{and}\qquad N>p^{\eta}$
by taking $r>1/\eta$.
The proof is quite short. From the Hölder inequality, it follows that
$|\mathfrak{B}(\bm{\alpha},\bm{\beta})|\leqslant\sum_{m\in{\mathcal{M}}}|\alpha_{m}|\left|\sum_{n\in{\mathcal{N}}}\beta_{n}\chi(m+n)\right|\leqslant\|\bm{\alpha}\|_{1}^{1-\frac{1}{r}}\|\bm{\alpha}\|_{2}^{\frac{1}{r}}W^{\frac{1}{2r}},$
where
$W=\sum_{m\in{\mathcal{M}}}\left|\sum_{n\in{\mathcal{N}}}\beta_{n}\chi(m+n)\right|^{2r}.$
The ingredient here is that one enlarges the sum over $m$ to ${\mathbb{F}}_{p}$, and ignores the structure of ${\mathcal{M}}$, so that
$W\leqslant\sum_{m\in{\mathbb{F}}_{p}}\left|\sum_{n\in{\mathcal{N}}}\beta_{n}\chi(m+n)\right|^{2r}.$
Now Theorem A.1 follows immediately from Lemma 3.1.
Acknowledgements
During the preparation of this work I.S. was supported in part by the Australian Research Council Grants DP230100530 and DP230100534, and P.X. was supported in part by the National Natural Science Foundation of China (No. 12025106).